### v2

# CHG Shapley: Efficient Data Valuation and Selection towards Trustworthy Machine Learning

Abstract
Understanding the decision-making process of machine learning models is crucial for ensuring trustworthy machine learning. Data Shapley, a landmark study on data valuation, advances this understanding by assessing the contribution of each datum to model accuracy. However, the resource-intensive and time-consuming nature of multiple model retraining poses challenges for applying Data Shapley to large datasets. To address this, we propose the CHG (Conduct of Hardness and Gradient) score, which approximates the utility of each data subset on model accuracy during a single model training. By deriving the closed-form expression of the Shapley value for each data point under the CHG score utility function, we reduce the computational complexity to the equivalent of a single model retraining, an exponential improvement over existing methods. Additionally, we employ CHG Shapley for real-time data selection, demonstrating its effectiveness in identifying high-value and noisy data. CHG Shapley facilitates trustworthy model training through efficient data valuation, introducing a novel data-centric perspective on trustworthy machine learning.
1 Introduction
The central problem of trustworthy machine learning is explaining the decision-making process of models to enhance the transparency of data-driven algorithms. However, the high complexity of machine learning model training and inference processes obscures an intuitive understanding of their internal mechanisms. Approaching trustworthy machine learning from a data-centric perspective (Liu et al., 2023) offers a new perspective for research. For trustworthy model inference, a representative algorithm is the SHAP (Lundberg and Lee, 2017), which quantitatively attributes model outputs to input features, clarifying which features influence specific results the most. SHAP and its variants(Kwon and Zou, 2022b) are widely applied in data analysis and healthcare. For trustworthy model training, the Data Shapley algorithm (Ghorbani and Zou, 2019) stands out. It quantitatively attributes a model’s performance to each training data point, identifying valuable data that improves performance and noisy data that degrades it. Data valuation not only informs us about the extent to which each training data impacts model performance but also underpins subsequent tasks such as data selection, data acquisition, data cleaning, etc., and supports the development of data markets (Mazumder et al., 2023).
The impressive effectiveness of both SHAP and Data Shapley algorithms is rooted in the Shapley value (Shapley, 1953). The unique feature of the Shapley value lies in its ability to accurately and fairly allocate contributions to each factor in decision-making processes where multiple factors interact with each other. However, the exact computation of the Shapley value is $\mathcal{O}(2^{n})$, where $n$ is the number of participants. Consequently, even when employing estimation methods akin to linear least squares regression (Lundberg and Lee, 2017) or Monte Carlo (Ghorbani and Zou, 2019), the efficiency of SHAP and Data Shapley algorithms sometimes fails to meet requirements. For example, the most efficient known data valuation method is KNNShapley (Jia et al., 2019) designed for the k-nearest neighbors algorithm, requiring $\mathcal{O}(\sqrt{n}\log(n)^{2})$ times of model retraining, where $n$ is the number of training data.
In this paper, we focus on the efficiency problem of data valuation on large-scale datasets. Instead of investigating more efficient and robust algorithms to approximate the Shapley value like in (Lundberg and Lee, 2017; Wang and Jia, 2023; Wang et al., 2024b), we concentrate on estimating the utility function. The underlying idea is that if the accuracy estimation function of a model trained on a subset can be analytically expressed through some features of the subset, then it becomes possible to directly derive an analytic expression for the Shapley value of each data point.
Our contributions are summarized as follows:
Efficient and Large-Scale Data Valuation Method: We introduce the CHG (Conduct of Hardness and Gradient) score, which assesses the influence of a data subset on model accuracy. By deriving the analytical expression for the Shapley value of each data point under this utility function, we significantly improve upon the $\mathcal{O}(n^{2}\log n)$ time complexity of Data Shapley. Using our proposed CHG Shapley method, valuing the full training dataset takes at most twice the time of conventional training, equivalent to one additional model retraining.
Real-Time Selection of Training Data in Large-Scale Datasets Based on Data Valuation: We employ CHG Shapley for real-time data selection. Experiments involving data-efficient training and model training under noise labels demonstrate the effectiveness of CHG Shapley in identifying high-value and noisy data.
A New Data-Centric Perspective on Trustworthy Machine Learning: CHG Shapley is a parameter-free method, deriving its benefits solely from a deeper understanding of the data. We believe that approaches rooted in data valuation have the potential to enhance understanding of model mechanisms, thereby advancing trustworthy machine learning.
2 Background and Related Works
Data-centric AI. For an extended period, the machine learning research community has predominantly concentrated on model development rather than on the underlying datasets (Mazumder et al., 2023). However, the inaccuracies, unfairness, and biases in models caused by data have become increasingly severe in real-world applications (Wang et al., 2023). Conducting trustworthy machine learning research from a data-centric perspective has garnered increasing attention from researchers (Liu et al., 2023). Research on data valuation can also be seen as an effort to understand the machine learning model training process from a data-centric perspective.
Shapley value. Supposing some players cooperate to finish a task, we denote the player set as $N$ and the utility function defined on every subset of $N$ is denoted as $U$, then the Shapley value (Shapley, 1953) is the unique solution concept that satisfies the following four axioms:
  • Dummy player: If $U(S\cup\{i\})=U(S)$ for all $S\subseteq N\setminus\{i\}$, then $\phi(i;U)=0$.
  • Symmetry: If $U(S\cup\{i\})=U(S\cup\{j\})$ for all $S\subseteq N\setminus\{i,j\}$, then $\phi(i;U)=\phi(j;U)$.
  • Linearity: For utility functions $U_{1},U_{2}$ and any $\alpha_{1},\alpha_{2}\in\mathbb{R}$, $\phi(i;\alpha_{1}U_{1}+\alpha_{2}U_{2})=\alpha_{1}\phi(i;U_{1})+\alpha_{2}\phi(i;U_{2})$.
  • Efficiency: $\sum_{i\in N}\phi(i;U)=U(N)$.
Definition 1 (Shapley (1953))
Given a player set $N$ and a utility function $U$, the Shapley value of a player $i\in N$ is defined as
$\phi_{i}\left(U\right):=\frac{1}{n}\sum_{k=1}^{n}{n-1\choose k-1}^{-1}\sum_{S\subseteq N\setminus\{i\},|S|=k-1}\left[U(S\cup i)-U(S)\right]$ (1)
Data valuation. Data valuation aims to quantitatively analyze the impact of training data on the performance of machine learning models, particularly deep neural networks. Determining the value of data can facilitate tasks such as data selection (Wang et al., 2024c), acquisition(Mazumder et al., 2023), and trading(Wang et al., 2024a). Since the performance of machine learning models is derived from the collective contribution of all data points, a fair data valuation algorithm is essential to accurately attribute each data point’s contribution to the model’s performance.
A milestone in data valuation research is the Data Shapley algorithm developed by Ghorbani and Zou (2019). Data Shapley builds upon the Shapley value formula by treating all data in the training dataset as players in a cooperative game of "training the model together". $U(S)$ represents the accuracy achieved by training the model on dataset subset $S$. Consequently, the Shapley value computed for each data point serves as an indication of its contribution to model performance. Data points with lower values are considered more likely to be harmful, while those with higher values are deemed to be lacking in the model’s representation. Removing low-value data points or acquiring more high-value data points can lead to improved model accuracy, a hypothesis supported by related experiments. Following Data Shapley, several subsequent have been proposed.
KNN Shapley(Jia et al., 2019) focus on the Shapley value of the k-nearest neighbors algorithm, proposing the closed-form expression of the Shapley value of the KNN scenario, notably improving computational efficiency to $\mathcal{O}(\sqrt{n}\log(n)^{2})$.
Beta Shapley (Kwon and Zou, 2022a) was developed to address an issue identified in Data Shapley regarding the uniform weight allocation for sets with different cardinalities. By relaxing the efficiency axiom in the Shapley value and utilizing a beta distribution to characterize the weights of sets of different cardinalities, Beta Shapley performs better than Data Shapley in tasks such as detecting mislabeled training data.
Data Banzhaf (Wang and Jia, 2023) aims to mitigate the noise introduced by stochastic gradient algorithms on model accuracy (i.e., utility function) during the model training process. Banzhaf value has also been demonstrated to be the most robust solution concept among all semi-values (i.e., Shapley values without Efficiency restrictions) in the presence of noise in the utility function. Therefore, Data Banzhaf is well-suited for evaluating training data in deep learning scenarios.
Additionally, there has been a surge in work on machine learning systems that utilize Shapley values to compute data values (Karlaš et al., 2024). All these studies utilize the Shapley value or its variants as their theoretical foundation because it is widely acknowledged in cooperative game theory as the fairest method for allocating contributions (Algaba et al., 2019).
To provide a more detailed assessment of various data valuation algorithms, Jiang et al. (2023) introduced the OpenDataVal benchmark library. This benchmark library abstracts away details such as model training, allowing researchers to focus solely on algorithm design for data valuation. We also use this benchmark to evaluate our proposed data valuation algorithm.
3 Efficient Data valuation
"What’s the value of your data?" Instead of directly assessing the contribution of each data point to model accuracy, we initially explore another pertinent question: "Which data holds greater value in the training process?" In the realm of data selection, numerous significant studies offer varying perspectives. Some focus on the utility of individual data, employing methods such as influence functions (Koh and Liang, 2017; Yang et al., 2023; Chhabra et al., 2024), EL2N score (Paul et al., 2021), InfoBatch (Qin et al., 2024), and worst-case training methods (Huang et al., 2022). These approaches utilize data information and model parameters to calculate a score (e.g., loss) for data selection. Conversely, another approach emphasizes the utility of data subsets (Killamsetty et al., 2021a, b; Zhao et al., 2021), aiming to identify representative subsets of the original dataset that can achieve comparable model accuracy, often utilizing corresponding gradient subsets. We reconcile these two perspectives and introduce the CHG score as a measure of data subset accuracy.
3.1 Preliminaries
Let $N=\{1,\dots,n\}$ denotes a labeled training set. The objective is to assign a scalar to each training data point based on its contribution to the model’s performance. A utility function $U:2^{N}\rightarrow R$, mapping from any subset $S$ of the training set (including the empty set) to the performance of the model trained on the subset (denoted as $U(S)$), is usually used for data valuation. In Data Shaply(Ghorbani and Zou, 2019), the accuracy of the model on a hold-out test set is used as $U(S)$.
3.2 CHG Score as a Utility Function for Subset Model Performance
The opacity of $U(S)$, or the fact that $U(S)$ can only be obtained after training a model on $S$, renders marginal contribution-based data valuation algorithms such as Data Shapley (Ghorbani and Zou, 2019), Beta Shapley (Kwon and Zou, 2022a), Data Banzhaf (Wang and Jia, 2023), AME (Lin et al., 2022), and LOO (leave-one-out error) time-consuming and computationally intensive.
On the contrary, if we can determine the model’s accuracy on $S$ without training on it, or even derive the expression for $U(S)$ in terms of $S$, then we can significantly reduce the model training time and expedite the calculation of each data’s value. To achieve this, we first investigate the impact of training the model using a subset of data $S$.
Here we focus on the model training process. Assuming that in the current iteration, the model parameters are denoted by $\theta$. We use $f(\theta):=\frac{1}{N}\sum_{i\in N}f(i;\theta)$ to represent the loss function of the neural network across the entire training dataset. A smaller loss indicates higher model accuracy on the training dataset.
Assuming at this point, no overfitting has occurred, i.e., the model’s accuracy on the training and test datasets is comparable, we can reasonably utilize the training dataset’s loss to estimate the model’s accuracy on the test dataset. If we update the parameters by training on the subset $S$ using gradient descent, then the next update direction of the parameters is $x:=\frac{1}{|S|}\sum_{i\in S}x_{i}$, where $x_{i}$ represents the gradient of the $i$-th data point. After one step of gradient descent, $\theta$ becomes $\theta-\eta x$. Subsequently, based on the following lemma, $\|\nabla f(\theta)\|_{2}^{2}-\|\nabla f(\theta)-x\|_{2}^{2}$ can be considered as a metric for evaluating the model’s performance.
Lemma 1 ((Nesterov, 2018))
Let $f$ be any differentiable function with $L$-Lipschitz continuous gradient, $\theta_{x}=\theta-\eta x$, and $\eta=1/L$ [^1] This is a commonly used assignment method, for example, in proving the convergence of gradient descent for non-convex functions to local minima (Nesterov, 2018), $\eta$ was set as $1/L$, then we have:
$f(\theta_{x})\leq f(\theta)+\langle\nabla f(\theta),\theta_{x}-\theta\rangle+\frac{L}{2}\|\theta_{x}-\theta\|_{2}^{2}=f(\theta)-\frac{\eta}{2}(\|\nabla f(\theta)||_{2}^{2}-\|\nabla f(\theta)-x||_{2}^{2})$ (2)
Proof: The first inequality is due to the definition of $L$-Lipschitz continuity.
When using traditional empirical risk minimization, measuring the model’s performance using the gradients of the training set on the training process is acceptable. However, inspired by the principles of worst-case training method (Huang et al., 2022) and InfoBatch (Qin et al., 2024), we introduce a method to incorporate the difficulty of data points into the model optimization process.
Specifically, our optimization objective is to minimize $\frac{1}{N}\sum_{i\in N}h_{i}f(i;\theta)$ rather than $\frac{1}{N}\sum_{i\in N}f(i;\theta)$ (i.e., empirical risk minimization), where $h_{i}$ is equal to $f(i;\theta)$ but does not participate in model backpropagation. $h_{i}$ can be viewed as the hardness of learning this data point. Generally, when there is no label noise, focusing on hard-to-learn data points can help improve the model’s generalization ability (Huang et al., 2022) and may reduce training time (Qin et al., 2024). Then the gradient of $i$-th data changes from $x_{i}$ to $h_{i}x_{i}$, and then we define the Conduct of Hardness and Gradient score (CHG score) of a data subset $S$ as $\|\frac{1}{N}\sum_{i\in N}l_{i}\nabla f(i;\theta)\|_{2}^{2}-\|\frac{1}{N}\sum_{i\in N}l_{i}\nabla f(i;\theta)-\frac{1}{S}\sum_{i\in S}l_{i}\nabla f(i;\theta)\|_{2}^{2}$. We use this score to measure the influence of data subsets on model accuracy. We also investigate the performance of the Hardness score ( $\frac{1}{S}\sum_{i\in S}l_{i}$) and Gradient Score ( $\|\frac{1}{N}\sum_{i\in N}\nabla f(i;\theta)\|_{2}^{2}-\|\frac{1}{N}\sum_{i\in N}\nabla f(i;\theta)-\frac{1}{S}\sum_{i\in S}\nabla f(i;\theta)\|_{2}^{2}$) in Table 2, see Hardness Shapley and Gradient Shapley.
Algorithms Underlying Method Model Retraining Complexity Quality of Data Values
Noisy Label Noisy Feature Point Point
Detection Detection Removal Addition
LOO Marginal contribution $n$ + +
Data Shapley (Ghorbani and Zou, 2019) Marginal contribution $\mathcal{O}(n^{2}\log n)$ + + ++ ++
KNN Shapley (Jia et al., 2019) Marginal contribution NA + + ++ ++
Beta Shapley (Kwon and Zou, 2022a) Marginal contribution $\mathcal{O}(n^{2}\log n)$ + + ++ ++
Data Banzhaf (Wang and Jia, 2023) Marginal contribution $\mathcal{O}(n\log n)$ + +
AME (Lin et al., 2022) Marginal contribution $\mathcal{O}(n^{2}\log n)$ ++ +
Influence Function (Koh and Liang, 2017) Gradient NA - - + +
LAVA (Just et al., 2023) Gradient NA - ++ + +
DVRL (Yoon et al., 2020) Importance weight $1$ + - + ++
Data-OOB (Kwon and Zou, 2023) Out-of-bag estimate NA ++ + - ++
CHG Shapley (this paper) Marginal contribution, $1$ ++ + ++ NA
Gradient and Loss
Table 1: A taxonomy of data valuation algorithms, mainly abstracted from (Jiang et al., 2023). The symbol “NA” means the method is not based on model training, then “No Answer”. The symbols ‘- / + / ++’ indicate that a corresponding data valuation method achieves a ‘similar or worse / better / much better’ performance than a random baseline, respectively.
3.3 CHG Shapley
After obtaining the utility function (CHG score), which approximates the influence of data subsets on model accuracy using gradient and loss information during model training, we can derive the analytical expression for the Shapley value of each data point under this utility function. This enables efficient computation of the value of each data point in large-scale datasets.
Theorem 1
Supposing for any subset $S\subseteq N$, $U(S)=\|\alpha\|^{2}_{2}-\|\frac{1}{S}\sum_{i\in S}x_{i}-\alpha\|^{2}_{2}$, then the Shapley value of j-th data point can be expressed as:
$\phi_{j}(U)=$ (3)
$\left(-\frac{\sum_{k=1}^{n}\frac{1}{k^{2}}}{n}+\frac{2\sum_{k=1}^{n}\frac{1}{k}-3\sum_{k=1}^{n}\frac{1}{k^{2}}+\frac{1}{n}}{n(n-1)}+2\frac{2\sum_{k=1}^{n}\frac{1}{k}-2\sum_{k=1}^{n}\frac{1}{k^{2}}-1+\frac{1}{n}}{n(n-1)(n-2)}\right)\|x_{j}\|^{2}_{2}$
$-2\frac{\sum_{k=1}^{n}\frac{1}{k}-\sum_{k=1}^{n}\frac{1}{k^{2}}-\frac{1}{n}+\frac{1}{n^{2}}}{(n-1)(n-2)}\langle\sum_{i\in N}x_{i},x_{j}\rangle+\frac{2\sum_{k=1}^{n}\frac{1}{k}-2\sum_{k=1}^{n}\frac{1}{k^{2}}-1+\frac{1}{n}}{n(n-1)(n-2)}\|\sum_{i\in N}x_{i}\|^{2}_{2}$
$+\left(\frac{\sum_{k=1}^{n}\frac{1}{k^{2}}-\frac{1}{n}}{n(n-1)}-\frac{2\sum_{k=2}^{n}\frac{1}{k}-2\sum_{k=2}^{n}\frac{1}{k^{2}}-1+\frac{1}{n}}{n(n-1)(n-2)}\right)(\sum_{i\in N}\|x_{i}\|^{2}_{2})$
$+2\frac{\sum_{k=1}^{n}\frac{1}{k}-\frac{1}{n}}{n-1}\langle x_{j},\alpha\rangle-2\frac{\sum_{k=1}^{n}\frac{1}{k}-1}{n(n-1)}\langle\sum_{i\in N}x_{i},\alpha\rangle$
Proof see Appendix B
We implemented our CHG Shapley data valuation method (see Algorithm 1) based on the code foundation of OpenDataVal (Jiang et al., 2023) and conducted experiments on the CIFAR-10 embedding dataset on a NVIDIA A40 server. The model used is pre-trained ResNet50 + logistic regression, and we report the results in Table 1.
For noisy label detection tasks and comparisons with other data evaluation methods, see Picture 2. For point removal tasks, see Picture 1.
The experimental results demonstrate that CHG Shapley data valuation method not only exhibits excellent data valuation and noise data detection capabilities, but also achieves remarkable operational efficiency. This is made possible by our research into the utility function that measures the impact of data subsets on model accuracy. In the next section, we will further show its data selection ability based on data valuation and noise label detection.
Algorithm 1 CHG Shapley-based Data Valuation Algorithm
0:  Training dataset ($N$), model architecture ($\theta_{0}$)
0:  Value of every training data
1:  Initialization: Model parameter and training hyperparameters initialization
2:  for $k$-th training epoch do
3:     Acquire loss and last-layer gradients of each data, i.e., $l_{i},\nabla f(i;\theta_{k})$
4:     Calculate Shapley value of every data point based on Equation 3, with $\alpha=\frac{1}{N}\sum_{i\in N}l_{i}\nabla f(i;\theta)$ and $x_{i}=l_{i}\nabla f(i;\theta_{k})$
5:     Update model parameters $\theta_{k}$ to $\theta_{k+1}$
6:  end for
7:  Return each data’s average Shapley value
4 Data Selection
In this section, we introduce various implementation strategies and practical techniques to enhance the scalability and efficiency of CHG Shapley. The experiments are conducted on a NVIDIA A40 server.
Last-Layer Gradients: Modern deep learning models possess a vast number of parameters, resulting in high-dimensional gradients. To mitigate this issue, we adopt a last-layer gradient approximation as in(Killamsetty et al., 2021a, b). This technique focuses solely on the gradients of the last layer for neural networks in CHG Shapley, significantly improving the speed of the algorithm and other baseline methods.
Per-Class Approximations: To enhance model performance, we adopt a per-class approach. This method involves caluclate the CHG Shapley separately for each class, which means considering only the data instances belonging to each class in each iteration. This approach not only reduces memory requirements but also decreases computational overhead, thereby improving the overall efficiency of the CHG Shapley-based data selection algorithm.
CHG Shapley-based Data Selection Algorithm: Here, we detail the CHG Shapley Data Selection Algorithm, integrating the aforementioned techniques to ensure efficient and scalable data valuation and selection.
Algorithm 2 CHG Shapley-based Data Selection Algorithm
0:  Training dataset ($N$), model architecture ($\theta_{0}$), selection interval $R=20$, the fraction (denoted as $a$) of data available for each epoch.
0:  Model parameters
1:  Initialization: Model parameter and training hyperparameters initialization, and let the training subset $S=\emptyset$.
2:  for $k$-th training epoch do
3:     if  $k\mod R==0$ then
4:        Let training subset $S=\emptyset$
5:        for $c$-th class do
6:           Acquire loss and last-layer gradients of each data in this class, i.e., $l_{i},\nabla f(i;\theta_{k})$
7:           Calculate Shapley value of every data point in this class (denoted as $N_{c}$) based on Equation 3, with $\alpha=\frac{1}{N_{c}}\sum_{i\in N_{c}}l_{i}\nabla f(i;\theta)$ and $x_{i}=l_{i}\nabla f(i;\theta_{k})$
8:           Add the top $aN_{c}$ data points with the highest Shapley values to the training subset $S$
9:        end for
10:        Calculate the Shapley values of each data point in $S$ with $\alpha=\frac{1}{S}\sum_{i\in S}l_{i}\nabla f(i;\theta)$ and $x_{i}=l_{i}\nabla f(i;\theta_{k})$, and apply minmax normalization to Shapley values to get the weights (denoted as $\{w_{S}\}$) of each data point in $S$
11:     end if
12:     Train model on $S$ using weighted learning with $\{w_{S}\}$, updating its parameters to $\theta_{k+1}$
13:  end for
Results: We report the accuracy and time consumption comparisons in Table 2 under the data pruning setting. CHG Shapley demonstrates particularly strong performance when the selection ratio is small, highlighting its ability to identify high-value data effectively. Additionally, the time consumption remains acceptable even when considering data selection processes, underscoring the high efficiency of CHG Shapley as a data valuation method. This efficiency, combined with its accuracy, verifies that CHG Shapley is a practical and powerful tool for data selection and valuation in large-scale machine learning tasks.
The results presented in Table 3, where 30% of the class labels were randomly replaced with other labels, provide additional evidence of CHG Shapley’s capability to detect mislabeled data. CHG Shapley demonstrates superior performance compared to other methods in this scenario, reinforcing its effectiveness in identifying and addressing label noise.
Table 2: The accuracy (%) and time (s) comparison. All methods are trained with ResNet-18.
Dataset CIFAR10 CIFAR100 Tiny ImageNet Fraction 0.05 0.1 0.3 0.5 0.7 0.05 0.1 0.3 0.5 0.7 0.05 0.1 0.3 0.5 0.7 Accuracy (%) Full 95.51 77.56 62.43 Random 65.90 77.36 89.93 93.11 93.87 22.58 36.51 62.27 69.70 73.04 17.13 27.70 45.17 52.37 57.89 AdaptiveRandom 83.26 88.78 93.88 94.77 94.66 48.52 60.58 72.45 74.97 75.70 37.76 48.50 58.49 60.30 62.18 Glister 85.44 89.92 91.26 91.93 93.04 48.25 64.63 72.72 74.00 75.14 37.17 48.29 56.77 59.13 61.50 GradMatch (PB) 85.17 90.17 93.93 94.42 94.83 50.42 63.53 73.79 75.78 76.20 38.88 49.97 59.79 61.56 62.19 CHG Shapley 87.15 90.88 93.65 94.51 94.58 53.97 64.46 72.54 74.38 75.71 38.92 49.00 58.44 60.68 61.80 Hardness Shapley 54.96 68.21 84.18 87.02 - 34.54 53.11 66.20 69.10 - - - - - - Gradient Shapley 84.87 89.95 93.54 94.08 - 51.15 59.73 70.69 73.88 - - - - - - Time (s) Full 7745 8082 49747 Random 776 1140 1906 6489 8708 805 1645 3839 3170 5859 2742 5184 15323 24475 33910 AdaptiveRandom 502 1522 4037 6815 6574 746 1358 3324 4786 7166 3451 5852 21366 27659 37583 Glister 1600 2736 4713 11853 15740 1247 1885 4778 11762 14611 4382 7693 19210 28289 36558 GradMatch (PB) 1308 2129 4897 7661 9916 804 2192 4077 6095 9185 4455 6904 24238 30814 51780 CHG Shapley 1383 1768 3746 7356 9526 1258 2037 4706 7012 9657 7077 9887 25682 30762 38177 Hardness Shapley 1191 1741 3606 4878 - 1233 1836 3828 5824 - - - - - - Gradient Shapley 1365 1933 2714 4017 - 1427 2025 2730 4032 - - - - - -
Table 3: Comparison of performance under noise label setting.
Dataset CIFAR10 CIFAR100 Fraction 0.05 0.1 0.3 0.5 0.7 0.05 0.1 0.3 0.5 0.7 AdaptiveRandom 57.18 63.21 71.84 75.38 77.41 24.12 32.69 45.36 52.18 53.00 Glister 61.17 70.39 73.73 75.64 76.59 18.64 29.22 45.80 49.78 52.45 GradMatch (PB) 58.38 69.67 72.59 75.41 76.90 25.90 33.84 47.71 52.35 54.29 CHG Shapley 67.43 77.32 85.33 88.83 88.90 35.75 46.41 59.59 61.92 63.36
5 Discussion and Future Work
In this paper, we define the Conduct of Hardness and Gradient score (CHG score) of a data subset $S$ as $\|\frac{1}{N}\sum_{i\in N}l_{i}\nabla f(i;\theta)\|_{2}^{2}-\|\frac{1}{N}\sum_{i\in N}l_{i}\nabla f(i;\theta)-\frac{1}{S}\sum_{i\in S}l_{i}\nabla f(i;\theta)\|_{2}^{2}$, and use this score based on Euclidean distance to measure the influence of data subsets on model accuracy. However, it’s worth noting that many studies on data selection based on gradients utilize different distance measures between $\frac{1}{N}\sum_{i\in N}l_{i}\nabla f(i;\theta)$ and $\frac{1}{S}\sum_{i\in S}l_{i}\nabla f(i;\theta)$, such as cosine similarity (Zhao et al., 2021), inner product, L1 norm, or their combinations, and then optimize this metric to select data. However, whether these metrics can be used for data valuation remains unknown because, unlike the theoretical support for the aforementioned Euclidean distance, these metrics lack theoretical grounding. Whether cosine similarity, inner product, L1 norm, or their combinations can achieve this goal remains to be investigated.
On the other hand, if cosine similarity and L1 norm are used in our framework, it becomes challenging to derive the corresponding Shapley values. In other words, if the Euclidean distance cannot effectively measure the utility of dataset subsets for model accuracy, the entire data valuation approach proposed in this study becomes computationally infeasible. The root of this problem lies in the complexity of computing Shapley values, which is generally challenging and only tractable when the utility function follows a specific form (such as linear models, Euclidean distance, or inner product).
Appendix A Data Valuation results
Figure 1: Point removal experiment on CIFAR10-embeddings dataset.
Figure 2: Discover noisy datas experiment on CIFAR10-embeddings dataset.
Appendix B Proof of Theorem  1
From $U(S)=||\alpha||^{2}-||\frac{1}{S}\sum_{i\in S}x_{i}-\alpha||^{2}$ we have $U(S)=-||\frac{1}{S}\sum_{i\in S}x_{i}||^{2}+2\frac{1}{S}\sum_{i\in S}x_{i}\alpha=U_{1}(S)+U_{2}(S)$, where $U_{1}(S)=-||\frac{1}{S}\sum_{i\in S}x_{i}||^{2}$, $U_{2}(S)=2\frac{1}{S}\sum_{i\in S}x_{i}\alpha$.
For $U_{2}$, we have: $\phi_{j}(U_{2})=\left(\frac{1}{n}\sum_{k=1}^{n}\frac{1}{k}\right)x_{j}\alpha-\left(\frac{1}{n}\sum_{k=2}^{n}\frac{1}{k}\right)(\frac{1}{n-1}\sum_{i\in N/{j}}x_{i}\alpha)$
Then we have: $\phi_{j}(U_{2})=\left(\frac{1}{n-1}\sum_{k=1}^{n-1}\frac{1}{k}\right)x_{j}\alpha-\left(\frac{1}{n-1}\sum_{k=2}^{n}\frac{1}{k}\right)(\frac{1}{n}\sum_{i\in N}x_{i}\alpha)$
Then:
$U_{1}(S\cup j)-U_{1}(S)=-\frac{1}{(S+1)^{2}}x_{j}^{2}-\frac{2}{(S+1)^{2}}x_{j}(\sum_{i\in S}x_{i})+\frac{2S+1}{S^{2}(S+1)^{2}}(\sum_{i\in S}x_{i})^{2}.$
Because of the linearity of the Shapley Value, we can calculate the Shapley value under utility functions $U_{1}$ and $U_{2}$ separately, and $\phi_{j}(U)=\phi_{j}(U_{1})+\phi_{j}(U_{2})$.
$\phi_{j}(U_{1})$ $=-\left(\frac{1}{n}\sum_{k=1}^{n}\frac{1}{k^{2}}\right)x_{j}^{2}-\left(\frac{2}{n(n-1)}\sum_{k=2}^{n}\frac{(k-1)}{k^{2}}\right)(\sum_{i\in N/{j}}x_{i})x_{j}$
$+\left(\frac{1}{n(n-1)}\sum_{k=2}^{n}\frac{(2k-1)}{k^{2}(k-1)}\right)(\sum_{i\in N/{j}}x_{i}^{2})$
$+\left(\frac{1}{n(n-1)(n-2)}\sum_{k=3}^{n}\frac{(2k-1)(k-2)}{k^{2}(k-1)}\right)(\sum_{a\in N/{j}}\sum_{b\in N/{j},a\neq b}x_{a}x_{b})$
Then we have:
$\phi_{j}(U_{1})$ $=-\left(\frac{1}{n}\sum_{k=1}^{n}\frac{1}{k^{2}}\right)x_{j}^{2}+\left(\frac{2}{n(n-1)}\sum_{k=2}^{n}\frac{(k-1)}{k^{2}}\right)x_{j}^{2}-\left(\frac{2}{n(n-1)}\sum_{k=2}^{n}\frac{(k-1)}{k^{2}}\right)(\sum_{i\in N}x_{i})x_{j}$
$+\left(\frac{1}{n(n-1)}\sum_{k=2}^{n}\frac{(2k-1)}{k^{2}(k-1)}\right)(\sum_{i\in N}x_{i}^{2})-\left(\frac{1}{n(n-1)}\sum_{k=2}^{n}\frac{(2k-1)}{k^{2}(k-1)}\right)x_{j}^{2}$
$-\left(\frac{1}{n(n-1)(n-2)}\sum_{k=3}^{n}\frac{(2k-1)(k-2)}{k^{2}(k-1)}\right)(\sum_{i\in N}x_{i}^{2})$
$+\left(\frac{1}{n(n-1)(n-2)}\sum_{k=3}^{n}\frac{(2k-1)(k-2)}{k^{2}(k-1)}\right)(\sum_{a\in N}\sum_{b\in N}x_{a}x_{b})$
$-2\left(\frac{1}{n(n-1)(n-2)}\sum_{k=3}^{n}\frac{(2k-1)(k-2)}{k^{2}(k-1)}\right)(\sum_{a\in N}x_{a}x_{j})$
$+2\left(\frac{1}{n(n-1)(n-2)}\sum_{k=3}^{n}\frac{(2k-1)(k-2)}{k^{2}(k-1)}\right)x_{j}^{2}$
Then we have:
$\phi_{j}(U_{1})$
$=[-\frac{1}{n}\sum_{k=1}^{n}\frac{1}{k^{2}}+\frac{2}{n(n-1)}\sum_{k=2}^{n}\frac{(k-1)}{k^{2}}-\frac{1}{n(n-1)}\sum_{k=2}^{n}\frac{(2k-1)}{k^{2}(k-1)}+\frac{2}{n(n-1)(n-2)}\sum_{k=3}^{n}\frac{(2k-1)(k-2)}{k^{2}(k-1)}]x_{j}^{2}$
$-\left(\frac{2}{n(n-1)}\sum_{k=2}^{n}\frac{(k-1)}{k^{2}}\right)(\sum_{i\in N}x_{i})x_{j}-2\left(\frac{1}{n(n-1)(n-2)}\sum_{k=3}^{n}\frac{(2k-1)(k-2)}{k^{2}(k-1)}\right)(\sum_{i\in N}x_{i})x_{j}$
$+\left(\frac{1}{n(n-1)(n-2)}\sum_{k=3}^{n}\frac{(2k-1)(k-2)}{k^{2}(k-1)}\right)(\sum_{a\in N}\sum_{b\in N}x_{a}x_{b})$
$+\left(\frac{1}{n(n-1)}\sum_{k=2}^{n}\frac{(2k-1)}{k^{2}(k-1)}-\frac{1}{n(n-1)(n-2)}\sum_{k=3}^{n}\frac{(2k-1)(k-2)}{k^{2}(k-1)}\right)(\sum_{i\in N}x_{i}^{2})$
Then we have:
$\phi_{j}(U_{1})$
$=[-\frac{1}{n}\sum_{k=1}^{n}\frac{1}{k^{2}}+\frac{1}{n(n-1)}\left(2\sum_{k=1}^{n}\frac{1}{k}-3\sum_{k=1}^{n}\frac{1}{k^{2}}+\frac{1}{n}\right)+2\frac{2\sum_{k=2}^{n}\frac{1}{k}-2\sum_{k=2}^{n}\frac{1}{k^{2}}-1+\frac{1}{n}}{n(n-1)(n-2)}]x_{j}^{2}$
$-\frac{2}{n(n-1)}\left(\sum_{k=2}^{n}\frac{1}{k}-\sum_{k=2}^{n}\frac{1}{k^{2}}+\frac{1}{(n-2)}\left(2\sum_{k=2}^{n}\frac{1}{k}-2\sum_{k=2}^{n}\frac{1}{k^{2}}-1+\frac{1}{n}\right)\right)(\sum_{i\in N}x_{i})x_{j}$
$+\frac{1}{n(n-1)(n-2)}\left(2\sum_{k=2}^{n}\frac{1}{k}-2\sum_{k=2}^{n}\frac{1}{k^{2}}-1+\frac{1}{n}\right)(\sum_{a\in N}\sum_{b\in N}x_{a}x_{b})$
$+\left(\frac{1}{n(n-1)}\left(\sum_{k=2}^{n}\frac{1}{k^{2}}+1-\frac{1}{n}\right)-\frac{2\sum_{k=2}^{n}\frac{1}{k}-2\sum_{k=2}^{n}\frac{1}{k^{2}}-1+\frac{1}{n}}{n(n-1)(n-2)}\right)(\sum_{i\in N}x_{i}^{2})$
Then we have:
$\phi_{j}(U_{1})$
$=[-\frac{1}{n}\sum_{k=1}^{n}\frac{1}{k^{2}}+\frac{1}{n(n-1)}\left(2\sum_{k=1}^{n}\frac{1}{k}-3\sum_{k=1}^{n}\frac{1}{k^{2}}+\frac{1}{n}\right)+2\frac{2\sum_{k=1}^{n}\frac{1}{k}-2\sum_{k=1}^{n}\frac{1}{k^{2}}-1+\frac{1}{n}}{n(n-1)(n-2)}]x_{j}^{2}$
$-\frac{2}{(n-1)(n-2)}\left(\sum_{k=1}^{n}\frac{1}{k}-\sum_{k=1}^{n}\frac{1}{k^{2}}-\frac{(n-1)}{n^{2}}\right)(\sum_{i\in N}x_{i})x_{j}$
$+\frac{1}{n(n-1)(n-2)}\left(2\sum_{k=1}^{n}\frac{1}{k}-2\sum_{k=1}^{n}\frac{1}{k^{2}}-1+\frac{1}{n}\right)(\sum_{a\in N}\sum_{b\in N}x_{a}x_{b})$
$+\left(\frac{1}{n(n-1)}\left(\sum_{k=1}^{n}\frac{1}{k^{2}}-\frac{1}{n}\right)-\frac{2\sum_{k=2}^{n}\frac{1}{k}-2\sum_{k=2}^{n}\frac{1}{k^{2}}-1+\frac{1}{n}}{n(n-1)(n-2)}\right)(\sum_{i\in N}x_{i}^{2})$
so :
$\phi_{j}(U)$
$=[-\frac{1}{n}\sum_{k=1}^{n}\frac{1}{k^{2}}+\frac{1}{n(n-1)}\left(2\sum_{k=1}^{n}\frac{1}{k}-3\sum_{k=1}^{n}\frac{1}{k^{2}}+\frac{1}{n}\right)+2\frac{2\sum_{k=1}^{n}\frac{1}{k}-2\sum_{k=1}^{n}\frac{1}{k^{2}}-1+\frac{1}{n}}{n(n-1)(n-2)}]x_{j}^{2}$
$-\frac{2}{(n-1)(n-2)}\left(\sum_{k=1}^{n}\frac{1}{k}-\sum_{k=1}^{n}\frac{1}{k^{2}}-\frac{1}{n}+\frac{1}{n^{2}}\right)(\sum_{i\in N}x_{i})x_{j}$
$+\frac{1}{n(n-1)(n-2)}\left(2\sum_{k=1}^{n}\frac{1}{k}-2\sum_{k=1}^{n}\frac{1}{k^{2}}-1+\frac{1}{n}\right)(\sum_{a\in N}\sum_{b\in N}x_{a}x_{b})$
$+\left(\frac{1}{n(n-1)}\left(\sum_{k=1}^{n}\frac{1}{k^{2}}-\frac{1}{n}\right)-\frac{2\sum_{k=2}^{n}\frac{1}{k}-2\sum_{k=2}^{n}\frac{1}{k^{2}}-1+\frac{1}{n}}{n(n-1)(n-2)}\right)(\sum_{i\in N}x_{i}^{2})$
$+\frac{2}{n-1}\left(\sum_{k=1}^{n}\frac{1}{k}-\frac{1}{n}\right)x_{j}\alpha-\frac{2}{n(n-1)}\left(\sum_{k=1}^{n}\frac{1}{k}-1\right)(\sum_{i\in N}x_{i})\alpha$

### v3

@@
 Understanding the decision-making process of machine learning models is crucial for ensuring trustworthy machine learning.
-Data Shapley, a landmark study on data valuation, advances this understanding by assessing the contribution of each datum to model accuracy.
+Data Shapley, a landmark study on data valuation, advances this understanding by assessing the contribution of each datum to model performance.
 However, the resource-intensive and time-consuming nature of multiple model retraining poses challenges for applying Data Shapley to large datasets.
-To address this, we propose the CHG (Conduct of Hardness and Gradient) score, which approximates the utility of each data subset on model accuracy during a single model training.
-By deriving the closed-form expression of the Shapley value for each data point under the CHG score utility function, we reduce the computational complexity to the equivalent of a single model retraining, an exponential improvement over existing methods.
-Additionally, we employ CHG Shapley for real-time data selection, demonstrating its effectiveness in identifying high-value and noisy data.
-CHG Shapley facilitates trustworthy model training through efficient data valuation, introducing a novel data-centric perspective on trustworthy machine learning.
+To address this, we propose the CHG (compound of Hardness and Gradient) utility function, which approximates the utility of each data subset on model performance in every training epoch.
+By deriving the closed-form Shapley value for each data point using the CHG utility function, we reduce the computational complexity to that of a single model retraining, achieving a quadratic improvement over existing marginal contribution-based methods.
+We further leverage CHG Shapley for real-time data selection, conducting experiments across three settings: standard datasets, label noise datasets, and class imbalance datasets.
+These experiments demonstrate its effectiveness in identifying high-value and noisy data.
+By enabling efficient data valuation, CHG Shapley promotes trustworthy model training through a novel data-centric perspective.
+Our codes are available at https://github.com/caihuaiguang/CHG-Shapley-for-Data-Valuation and https://github.com/caihuaiguang/CHG-Shapley-for-Data-Selection.
 1 Introduction
@@
 Approaching trustworthy machine learning from a data-centric perspective (Liu et al., 2023) offers a new perspective for research.
-For trustworthy model inference, a representative algorithm is the SHAP (Lundberg and Lee, 2017), which quantitatively attributes model outputs to input features, clarifying which features influence specific results the most.
-SHAP and its variants(Kwon and Zou, 2022b) are widely applied in data analysis and healthcare.
-For trustworthy model training, the Data Shapley algorithm (Ghorbani and Zou, 2019) stands out.
+For trustworthy model inference, a representative algorithm is the SHAP (Lundberg & Lee, 2017), which quantitatively attributes model outputs to input features, clarifying which features influence specific results the most.
+SHAP and its variants(Kwon & Zou, 2022b) are widely applied in data analysis and healthcare.
+For trustworthy model training, the Data Shapley algorithm (Ghorbani & Zou, 2019) stands out.
 It quantitatively attributes a model’s performance to each training data point, identifying valuable data that improves performance and noisy data that degrades it.
-Data valuation not only informs us about the extent to which each training data impacts model performance but also underpins subsequent tasks such as data selection, data acquisition, data cleaning, etc., and supports the development of data markets (Mazumder et al., 2023).
+Data valuation reveals how much each training sample affects model performance, serving as a foundation for tasks like data selection, acquisition, and cleaning, while also facilitating the creation of data markets (Mazumder et al., 2023).
 The impressive effectiveness of both SHAP and Data Shapley algorithms is rooted in the Shapley value (Shapley, 1953).
 The unique feature of the Shapley value lies in its ability to accurately and fairly allocate contributions to each factor in decision-making processes where multiple factors interact with each other.
-However, the exact computation of the Shapley value is $\mathcal{O}(2^{n})$, where $n$ is the number of participants.
-Consequently, even when employing estimation methods akin to linear least squares regression (Lundberg and Lee, 2017) or Monte Carlo (Ghorbani and Zou, 2019), the efficiency of SHAP and Data Shapley algorithms sometimes fails to meet requirements.
-For example, the most efficient known data valuation method is KNNShapley (Jia et al., 2019) designed for the k-nearest neighbors algorithm, requiring $\mathcal{O}(\sqrt{n}\log(n)^{2})$ times of model retraining, where $n$ is the number of training data.
+However, the exact computation of the Shapley value is $\mathcal{O}(2^{n})$, where $n$ is the number of factors.
+Even with estimation techniques such as linear least squares regression (Lundberg & Lee, 2017) or Monte Carlo methods (Ghorbani & Zou, 2019), the efficiency of SHAP and Data Shapley algorithms struggles to scale with high-dimensional inputs or large datasets.
 In this paper, we focus on the efficiency problem of data valuation on large-scale datasets.
-Instead of investigating more efficient and robust algorithms to approximate the Shapley value like in (Lundberg and Lee, 2017; Wang and Jia, 2023; Wang et al., 2024b), we concentrate on estimating the utility function.
-The underlying idea is that if the accuracy estimation function of a model trained on a subset can be analytically expressed through some features of the subset, then it becomes possible to directly derive an analytic expression for the Shapley value of each data point.
+Instead of investigating more efficient and robust algorithms to approximate the Shapley value as in (Lundberg & Lee, 2017; Wang et al., 2024b), we concentrate on estimating the utility function.
+The key intuition is that if the performance of a model trained on a data subset can be expressed analytically, a closed-form solution for each datum’s Shapley value could be derived, offering a more computationally efficient solution.
 Our contributions are summarized as follows:
-Efficient and Large-Scale Data Valuation Method: We introduce the CHG (Conduct of Hardness and Gradient) score, which assesses the influence of a data subset on model accuracy.
-By deriving the analytical expression for the Shapley value of each data point under this utility function, we significantly improve upon the $\mathcal{O}(n^{2}\log n)$ time complexity of Data Shapley.
-Using our proposed CHG Shapley method, valuing the full training dataset takes at most twice the time of conventional training, equivalent to one additional model retraining.
-Real-Time Selection of Training Data in Large-Scale Datasets Based on Data Valuation: We employ CHG Shapley for real-time data selection.
-Experiments involving data-efficient training and model training under noise labels demonstrate the effectiveness of CHG Shapley in identifying high-value and noisy data.
-A New Data-Centric Perspective on Trustworthy Machine Learning: CHG Shapley is a parameter-free method, deriving its benefits solely from a deeper understanding of the data.
-We believe that approaches rooted in data valuation have the potential to enhance understanding of model mechanisms, thereby advancing trustworthy machine learning.
+Efficient and Large-Scale Data Valuation Method: We introduce the CHG (Compound of Hardness and Gradient) score, which assesses the influence of a data subset on model accuracy.
+By deriving the analytical expression for the Shapley value of each data point under this utility function, we significantly improve upon the $\mathcal{O}(n^{2}\log n)$ time complexity of Data Shapley to a single model training run.
+Real-Time Training Data Selection in Large Datasets: Calculating data value was considered computationally intensive, making real-time data selection based on data value impractical.
+Due to the efficient computation of CHG Shapley, we further employ it for real-time data selection.
+Experiments conducted across three settings—standard datasets, label noise datasets, and class imbalance datasets—demonstrate CHG Shapley’s effectiveness in identifying high-value and noisy data.
+A New Data-Centric Perspective on Trustworthy Machine Learning: In real-world scenarios, data may not only be subject to noise and class imbalance, but it can also involve more complex mixtures of these issues.
+CHG Shapley is a parameter-free method that derives its advantages solely from a deeper understanding of the data.
+We believe that the proposed method, grounded in data valuation, has the potential to enhance trustworthy model training in complex data distributions.
 2 Background and Related Works
@@
 Shapley value.
-Supposing some players cooperate to finish a task, we denote the player set as $N$ and the utility function defined on every subset of $N$ is denoted as $U$, then the Shapley value (Shapley, 1953) is the unique solution concept that satisfies the following four axioms:
+The Shapley value (Shapley, 1953) is widely regarded as the fairest method for allocating contributions (Rozemberczki et al., 2022; Algaba et al., 2019).
+Given a set of players $N$ and a utility function $U$ defined on subsets of $N$, the Shapley value is the only solution that satisfies the axioms of dummy player, symmetry, efficiency, and linearity:
   • Dummy player: If $U(S\cup\{i\})=U(S)$ for all $S\subseteq N\setminus\{i\}$, then $\phi(i;U)=0$.
@@
   • Linearity: For utility functions $U_{1},U_{2}$ and any $\alpha_{1},\alpha_{2}\in\mathbb{R}$, $\phi(i;\alpha_{1}U_{1}+\alpha_{2}U_{2})=\alpha_{1}\phi(i;U_{1})+\alpha_{2}\phi(i;U_{2})$.
-  • Efficiency: $\sum_{i\in N}\phi(i;U)=U(N)$.
+  • Efficiency: $\sum_{i\in N}\phi(i;U)=U(N)-U(\emptyset)$.
 Definition 1 (Shapley (1953))
@@
 $\phi_{i}\left(U\right):=\frac{1}{n}\sum_{k=1}^{n}{n-1\choose k-1}^{-1}\sum_{S\subseteq N\setminus\{i\},|S|=k-1}\left[U(S\cup i)-U(S)\right]$ (1)
+A more intuitive form for the Shapley value is:
+$\phi_{i}(U)=\mathbb{E}_{\pi\sim\Pi}\left[U\left(S_{\pi}^{i}\cup\{i\}\right)-U\left(S_{\pi}^{i}\right)\right].$ (2)
+Here, $\pi\sim\Pi$ refers to a uniformly random permutation of the player set $N$, and $S_{\pi}^{i}$ denotes the set of players preceding player $i$ in the permutation $\pi$.
+This expression highlights that the Shapley value for player $i$ captures the expected marginal contribution of that player across all possible subsets.
+Fig.
+1 illustrates the process of computing Shapley values.
+Figure 1: Shapley value’s calculation.
 Data valuation.
@@
 Since the performance of machine learning models is derived from the collective contribution of all data points, a fair data valuation algorithm is essential to accurately attribute each data point’s contribution to the model’s performance.
-A milestone in data valuation research is the Data Shapley algorithm developed by Ghorbani and Zou (2019).
-Data Shapley builds upon the Shapley value formula by treating all data in the training dataset as players in a cooperative game of "training the model together".
+A milestone in data valuation research is the Data Shapley algorithm developed by Ghorbani & Zou (2019).
+Data Shapley builds upon the Shapley value formula by treating all data in the training dataset as players in a cooperative game of “training the model together”.
 $U(S)$ represents the accuracy achieved by training the model on dataset subset $S$.
 Consequently, the Shapley value computed for each data point serves as an indication of its contribution to model performance.
-Data points with lower values are considered more likely to be harmful, while those with higher values are deemed to be lacking in the model’s representation.
-Removing low-value data points or acquiring more high-value data points can lead to improved model accuracy, a hypothesis supported by related experiments.
+Data points with lower values are considered more likely to be harmful, while those with higher values are deemed to be lacking in the dataset.
+Removing low-value or acquiring more high-value data can improve accuracy.
 Following Data Shapley, several subsequent have been proposed.
-KNN Shapley(Jia et al., 2019) focus on the Shapley value of the k-nearest neighbors algorithm, proposing the closed-form expression of the Shapley value of the KNN scenario, notably improving computational efficiency to $\mathcal{O}(\sqrt{n}\log(n)^{2})$.
-Beta Shapley (Kwon and Zou, 2022a) was developed to address an issue identified in Data Shapley regarding the uniform weight allocation for sets with different cardinalities.
-By relaxing the efficiency axiom in the Shapley value and utilizing a beta distribution to characterize the weights of sets of different cardinalities, Beta Shapley performs better than Data Shapley in tasks such as detecting mislabeled training data.
-Data Banzhaf (Wang and Jia, 2023) aims to mitigate the noise introduced by stochastic gradient algorithms on model accuracy (i.e., utility function) during the model training process.
-Banzhaf value has also been demonstrated to be the most robust solution concept among all semi-values (i.e., Shapley values without Efficiency restrictions) in the presence of noise in the utility function.
-Therefore, Data Banzhaf is well-suited for evaluating training data in deep learning scenarios.
-Additionally, there has been a surge in work on machine learning systems that utilize Shapley values to compute data values (Karlaš et al., 2024).
-All these studies utilize the Shapley value or its variants as their theoretical foundation because it is widely acknowledged in cooperative game theory as the fairest method for allocating contributions (Algaba et al., 2019).
-To provide a more detailed assessment of various data valuation algorithms, Jiang et al.
-(2023) introduced the OpenDataVal benchmark library.
-This benchmark library abstracts away details such as model training, allowing researchers to focus solely on algorithm design for data valuation.
-We also use this benchmark to evaluate our proposed data valuation algorithm.
+KNN Shapley (Jia et al., 2019) focuses on the Shapley value for k-nearest neighbors, offering a closed-form solution that improves computational efficiency to $\mathcal{O}(\sqrt{n}\log(n)^{2})$.
+Beta Shapley (Kwon & Zou, 2022a) addresses the uniform weighting issue in Data Shapley by relaxing the efficiency axiom and using a beta distribution, improving performance in tasks like mislabeled data detection.
+Data Banzhaf (Wang & Jia, 2023) reduces noise in model training by utilizing the Banzhaf value, shown to be robust in noisy environments.
+LAVA (Just et al., 2023) developed a proxy for validation performance, enabling the evaluation of data in a manner that is agnostic to the learning algorithms.
+DVRL (Kwon & Zou, 2023) and Data-OOB (Yoon et al., 2020) utilize reinforcement learning and out-of-bag estimation, respectively, to assess data values.
+Several other works use gradient information to compute the value of training data.
+Xu et al.
+(2021) propose the cosine GradE Shapley value (CGSV) to measure each agent’s contribution in a collaborative machine learning scenario.
+CGSV approximates an agent’s Shapley value during the aggregation process by using the cosine similarity between the agent’s gradient and the aggregated gradient of all agents.
+Pruthi et al.
+(2020) introduced TracIn, a method that measures the influence of a training example on a test example by computing the inner product between their respective gradients.
+Finally, Jiang et al.
+(2023) introduced the OpenDataVal benchmark, which focuses on data valuation algorithm design by abstracting model training.
+We also use this benchmark to evaluate our proposed algorithm in data valuation experiments.
+Data selection.
+There are two main streams in the data selection field.
+The first focuses on determining the scores of individual data points, employing methods such as influence functions (Koh & Liang, 2017; Yang et al., 2023; Chhabra et al., 2024), the EL2N score (Paul et al., 2021), and loss (i.e., hardness) metrics (e.g., InfoBatch (Qin et al., 2024), worst-case training (Huang et al., 2022)).
+These approaches typically use the information from a single data point and model parameters to calculate scores for that point, subsequently selecting the data points with the highest scores.
+Recently, Xia et al.
+(2024) proposed LESS, which sets the gradient mean of data within a specific downstream task as the target gradient and employs the inner product or cosine similarity to select data with the most similar gradients from a large collection of instruction datasets for model fine-tuning.
+Their experiments (Xia et al., 2024) show that training on just 5% of the data selected by LESS often outperforms training on the full dataset across various downstream tasks.
+The second approach, including coreset methods such as GradMatch (Killamsetty et al., 2021a) and Glister (Killamsetty et al., 2021b), as well as data distillation (Zhao et al., 2021), emphasizes the utility of data subsets.
+These methods aim to identify a subset that effectively represents the original dataset while achieving comparable model accuracy.
+They typically leverage the gradients of individual training data points to form a subset whose gradient direction aligns with that of the validation set.
+Our data selection method, CHG Shapley, which is grounded in data valuation principles, represents a fusion of these two streams: it leverages the utility of different training subsets to calculate a score (i.e., Shapley value) for each individual data point.
+We will compare our method with these two approaches in data selection experiments, demonstrating its ability to identify high-value (i.e., representative) data and low-value (i.e., noisy) data.
 3 Efficient Data valuation
-"What’s the value of your data?" Instead of directly assessing the contribution of each data point to model accuracy, we initially explore another pertinent question: "Which data holds greater value in the training process?" In the realm of data selection, numerous significant studies offer varying perspectives.
-Some focus on the utility of individual data, employing methods such as influence functions (Koh and Liang, 2017; Yang et al., 2023; Chhabra et al., 2024), EL2N score (Paul et al., 2021), InfoBatch (Qin et al., 2024), and worst-case training methods (Huang et al., 2022).
-These approaches utilize data information and model parameters to calculate a score (e.g., loss) for data selection.
-Conversely, another approach emphasizes the utility of data subsets (Killamsetty et al., 2021a, b; Zhao et al., 2021), aiming to identify representative subsets of the original dataset that can achieve comparable model accuracy, often utilizing corresponding gradient subsets.
-We reconcile these two perspectives and introduce the CHG score as a measure of data subset accuracy.
 3.1 Preliminaries
-Let $N=\{1,\dots,n\}$ denotes a labeled training set.
-The objective is to assign a scalar to each training data point based on its contribution to the model’s performance.
-A utility function $U:2^{N}\rightarrow R$, mapping from any subset $S$ of the training set (including the empty set) to the performance of the model trained on the subset (denoted as $U(S)$), is usually used for data valuation.
-In Data Shaply(Ghorbani and Zou, 2019), the accuracy of the model on a hold-out test set is used as $U(S)$.
+Let $N=\{1,\dots,n\}$ denote a labeled training set.
+The objective is to assign a scalar value to each training data point based on its contribution to the model’s performance.
+A utility function $U:2^{N}\rightarrow\mathbb{R}$ maps any subset $S$ of the training set (including the empty set) to the performance of the model trained on that subset, denoted as $U(S)$.
+In Data Shapley (Ghorbani & Zou, 2019), the model’s accuracy on a hold-out test set is utilized as $U(S)$.
+Besides, similar to Data Shapley, we do not impose any distributional assumptions on the data, and independence among the data points is not required.
 3.2 CHG Score as a Utility Function for Subset Model Performance
-The opacity of $U(S)$, or the fact that $U(S)$ can only be obtained after training a model on $S$, renders marginal contribution-based data valuation algorithms such as Data Shapley (Ghorbani and Zou, 2019), Beta Shapley (Kwon and Zou, 2022a), Data Banzhaf (Wang and Jia, 2023), AME (Lin et al., 2022), and LOO (leave-one-out error) time-consuming and computationally intensive.
-On the contrary, if we can determine the model’s accuracy on $S$ without training on it, or even derive the expression for $U(S)$ in terms of $S$, then we can significantly reduce the model training time and expedite the calculation of each data’s value.
+The opacity of $U(S)$, or the fact that $U(S)$ can only be obtained after training a model on $S$, renders marginal contribution-based data valuation algorithms such as Data Shapley (Ghorbani & Zou, 2019), Beta Shapley (Kwon & Zou, 2022a), Data Banzhaf (Wang & Jia, 2023), AME (Lin et al., 2022), and LOO (leave-one-out error) time-consuming and computationally intensive.
+On the contrary, if we can determine the model’s performance on $S$ without training on it (Wu et al., 2022; Ki et al., 2023), or even derive the expression for $U(S)$ in terms of $S$, then we may significantly reduce the model training time and expedite the calculation of each data’s value.
 To achieve this, we first investigate the impact of training the model using a subset of data $S$.
-Here we focus on the model training process.
-Assuming that in the current iteration, the model parameters are denoted by $\theta$.
-We use $f(\theta):=\frac{1}{N}\sum_{i\in N}f(i;\theta)$ to represent the loss function of the neural network across the entire training dataset.
-A smaller loss indicates higher model accuracy on the training dataset.
-Assuming at this point, no overfitting has occurred, i.e., the model’s accuracy on the training and test datasets is comparable, we can reasonably utilize the training dataset’s loss to estimate the model’s accuracy on the test dataset.
-If we update the parameters by training on the subset $S$ using gradient descent, then the next update direction of the parameters is $x:=\frac{1}{|S|}\sum_{i\in S}x_{i}$, where $x_{i}$ represents the gradient of the $i$-th data point.
-After one step of gradient descent, $\theta$ becomes $\theta-\eta x$.
-Subsequently, based on the following lemma, $\|\nabla f(\theta)\|_{2}^{2}-\|\nabla f(\theta)-x\|_{2}^{2}$ can be considered as a metric for evaluating the model’s performance.
+The objective of model training is to minimize training loss while ensuring generalization (Zhang et al., 2017) through an appropriate dataset, sufficient data quantity, model complexity, and regularization techniques, etc.
+Once the dataset, model architecture, regularization techniques, and other methods that ensure generalization are established, the goal with all training data is to minimize training loss as much as possible.
+Therefore, in our study, the utility function is defined as the degree of decrease in training loss.
+Let the loss function of the neural network across the entire training dataset be represented as $f(\theta):=\frac{1}{N}\sum_{i\in N}f(i;\theta)$.
+When training the model on a subset $S$ using gradient descent, the next update direction for the parameters is given by $x:=\frac{1}{|S|}\sum_{i\in S}x_{i}$, where $x_{i}$ represents the gradient of the $i$-th data point.
+After one step of gradient descent, the parameters are updated as $\theta\leftarrow\theta-\eta x$.
+Based on the following lemma, the term $\|\nabla f(\theta)\|_{2}^{2}-\|\nabla f(\theta)-x\|_{2}^{2}$ can be employed as a utility function to evaluate the impacts of the subset $S$ on the training loss.
 Lemma 1 ((Nesterov, 2018))
-Let $f$ be any differentiable function with $L$-Lipschitz continuous gradient, $\theta_{x}=\theta-\eta x$, and $\eta=1/L$ [^1] This is a commonly used assignment method, for example, in proving the convergence of gradient descent for non-convex functions to local minima (Nesterov, 2018), $\eta$ was set as $1/L$, then we have:
-$f(\theta_{x})\leq f(\theta)+\langle\nabla f(\theta),\theta_{x}-\theta\rangle+\frac{L}{2}\|\theta_{x}-\theta\|_{2}^{2}=f(\theta)-\frac{\eta}{2}(\|\nabla f(\theta)||_{2}^{2}-\|\nabla f(\theta)-x||_{2}^{2})$ (2)
-Proof: The first inequality is due to the definition of $L$-Lipschitz continuity.
-When using traditional empirical risk minimization, measuring the model’s performance using the gradients of the training set on the training process is acceptable.
-However, inspired by the principles of worst-case training method (Huang et al., 2022) and InfoBatch (Qin et al., 2024), we introduce a method to incorporate the difficulty of data points into the model optimization process.
-Specifically, our optimization objective is to minimize $\frac{1}{N}\sum_{i\in N}h_{i}f(i;\theta)$ rather than $\frac{1}{N}\sum_{i\in N}f(i;\theta)$ (i.e., empirical risk minimization), where $h_{i}$ is equal to $f(i;\theta)$ but does not participate in model backpropagation.
-$h_{i}$ can be viewed as the hardness of learning this data point.
-Generally, when there is no label noise, focusing on hard-to-learn data points can help improve the model’s generalization ability (Huang et al., 2022) and may reduce training time (Qin et al., 2024).
-Then the gradient of $i$-th data changes from $x_{i}$ to $h_{i}x_{i}$, and then we define the Conduct of Hardness and Gradient score (CHG score) of a data subset $S$ as $\|\frac{1}{N}\sum_{i\in N}l_{i}\nabla f(i;\theta)\|_{2}^{2}-\|\frac{1}{N}\sum_{i\in N}l_{i}\nabla f(i;\theta)-\frac{1}{S}\sum_{i\in S}l_{i}\nabla f(i;\theta)\|_{2}^{2}$.
-We use this score to measure the influence of data subsets on model accuracy.
-We also investigate the performance of the Hardness score ( $\frac{1}{S}\sum_{i\in S}l_{i}$) and Gradient Score ( $\|\frac{1}{N}\sum_{i\in N}\nabla f(i;\theta)\|_{2}^{2}-\|\frac{1}{N}\sum_{i\in N}\nabla f(i;\theta)-\frac{1}{S}\sum_{i\in S}\nabla f(i;\theta)\|_{2}^{2}$) in Table 2, see Hardness Shapley and Gradient Shapley.
+Let $f$ be any differentiable function with $L$-Lipschitz continuous gradient, $\theta_{x}=\theta-\eta x$, and $\eta=1/L$ [^1] This is a commonly used assignment method, for example, in proving the convergence of gradient descent for non-convex functions to local minima (Nesterov, 2018), $\eta$ was set as $1/L$., then we have:
+$f(\theta_{x})\leq f(\theta)-\frac{1}{2L}(\|\nabla f(\theta)||_{2}^{2}-\|\nabla f(\theta)-x\|_{2}^{2})$ (3)
+Proof: Due to the definition of $L$-Lipschitz continuity, We have $f(\theta_{x})\leq f(\theta)+\langle\nabla f(\theta),\theta_{x}-\theta\rangle+\frac{L}{2}\|\theta_{x}-\theta\|_{2}^{2}$ Plugging in $\theta_{x}=\theta-\eta x$ to it, we get $\langle\nabla f(\theta),\theta_{x}-\theta\rangle=\langle\nabla f(\theta),-\eta x\rangle$ and $\|\theta_{x}-\theta\|_{2}^{2}=\eta^{2}\|x\|_{2}^{2}$.
+And with $\eta=1/L$, we then have $f(\theta)+\langle\nabla f(\theta),\theta_{x}-\theta\rangle+\frac{L}{2}\|\theta_{x}-\theta\|_{2}^{2}=f(\theta)+\langle\nabla f(\theta),-\eta x\rangle+\frac{L}{2}\eta^{2}\|x\|_{2}^{2}=f(\theta)-\frac{1}{2L}(2\langle\nabla f(\theta),x\rangle-\|x\|_{2}^{2})=f(\theta)-\frac{1}{2L}(\|\nabla f(\theta)||_{2}^{2}-\|\nabla f(\theta)-x\|_{2}^{2})$.
+Lemma 1 is particularly important to our whole method.
+It suggests that the Euclidean distance of two gradients is not just a similarity metric like inner product or cosine, but also gives an upper bound of the training loss, which then serves as the utility function of any data subset to calculate the Shapley value of each single data.
+In general, when there is no label noise, focusing on hard-to-learn data points can enhance the model’s generalization ability (Huang et al., 2022) and reduce training time without affecting test performance (Qin et al., 2024).
+Thus, we propose to incorporate the hardness of data points into the model optimization process when there is little to no label noise.
+Specifically, our optimization objective is to minimize $\frac{1}{N}\sum_{i\in N}h_{i}f(i;\theta)$ rather than usual empirical risk $\frac{1}{N}\sum_{i\in N}f(i;\theta)$, where $h_{i}$ is equal to $f(i;\theta)$ but does not participate in model backpropagation.
+Here, $h_{i}$ can also be interpreted as an unnormalized probability of selecting data point $i$ based on its hardness.
+Then the gradient of $i$-th data changes from $x_{i}$ to $h_{i}x_{i}$, and then we define the Compound of Hardness and Gradient score (CHG score) of a data subset $S$ as $U(S):=\|\frac{1}{N}\sum_{i\in N}h_{i}\nabla f(i;\theta)\|_{2}^{2}-\|\frac{1}{N}\sum_{i\in N}h_{i}\nabla f(i;\theta)-\frac{1}{S}\sum_{i\in S}h_{i}\nabla f(i;\theta)\|_{2}^{2}$.
+When the subset $S=\emptyset$, the produced utility $U(S)$ is 0.
+However, when label noise cannot be ignored, the hardness metric becomes unreliable, as a well-performing classifier may assign a high hardness value to a training data point with an incorrect label.
+In this case, we still utilize the original score $\|\frac{1}{N}\sum_{i\in N}\nabla f(i;\theta)\|_{2}^{2}-\|\frac{1}{N}\sum_{i\in N}\nabla f(i;\theta)-\frac{1}{S}\sum_{i\in S}\nabla f(i;\theta)\|_{2}^{2},$ the corresponding method for calculating the Shapley value based on Equation 4 is referred to as GradE Shapley (Gradient Euclidean Shapley).
+We also denote the direct use of the hardness $h_{i}$ as the Shapley value of data point $i$ as Hardness Shapley.
+We compared these three methods across various dataset settings in our data selection experiments.
 Algorithms Underlying Method Model Retraining Complexity Quality of Data Values
@@
 LOO Marginal contribution $n$ + +
-Data Shapley (Ghorbani and Zou, 2019) Marginal contribution $\mathcal{O}(n^{2}\log n)$ + + ++ ++
+Data Shapley (Ghorbani & Zou, 2019) Marginal contribution $\mathcal{O}(n^{2}\log n)$ + + ++ ++
 KNN Shapley (Jia et al., 2019) Marginal contribution NA + + ++ ++
-Beta Shapley (Kwon and Zou, 2022a) Marginal contribution $\mathcal{O}(n^{2}\log n)$ + + ++ ++
-Data Banzhaf (Wang and Jia, 2023) Marginal contribution $\mathcal{O}(n\log n)$ + +
+Beta Shapley (Kwon & Zou, 2022a) Marginal contribution $\mathcal{O}(n^{2}\log n)$ + + ++ ++
+Data Banzhaf (Wang & Jia, 2023) Marginal contribution $\mathcal{O}(n\log n)$ + +
 AME (Lin et al., 2022) Marginal contribution $\mathcal{O}(n^{2}\log n)$ ++ +
-Influence Function (Koh and Liang, 2017) Gradient NA - - + +
+Influence Function (Koh & Liang, 2017) Gradient NA - - + +
 LAVA (Just et al., 2023) Gradient NA - ++ + +
 DVRL (Yoon et al., 2020) Importance weight $1$ + - + ++
-Data-OOB (Kwon and Zou, 2023) Out-of-bag estimate NA ++ + - ++
-CHG Shapley (this paper) Marginal contribution, $1$ ++ + ++ NA
+Data-OOB (Kwon & Zou, 2023) Out-of-bag estimate NA ++ + - ++
+CHG Shapley (this paper) Marginal contribution, $1$ ++ ++ ++ NA
 Gradient and Loss
@@
 The symbols ‘- / + / ++’ indicate that a corresponding data valuation method achieves a ‘similar or worse / better / much better’ performance than a random baseline, respectively.
-3.3 CHG Shapley
-After obtaining the utility function (CHG score), which approximates the influence of data subsets on model accuracy using gradient and loss information during model training, we can derive the analytical expression for the Shapley value of each data point under this utility function.
-This enables efficient computation of the value of each data point in large-scale datasets.
+Detailed experimental results comparing these data valuation methods are provided in Appendix A.
+3.3 CHG Shapley-based Data Valuation Algorithm
+After obtaining the CHG score, which approximates the influence of data subsets on training loss using gradient and loss information during model training, we can derive the analytical expression for the Shapley value of each data point under this utility function.
+This method, denoted as CHG Shapley, enables efficient computation of the value of each data point in large-scale datasets.
 Theorem 1
@@
 $+2\frac{\sum_{k=1}^{n}\frac{1}{k}-\frac{1}{n}}{n-1}\langle x_{j},\alpha\rangle-2\frac{\sum_{k=1}^{n}\frac{1}{k}-1}{n(n-1)}\langle\sum_{i\in N}x_{i},\alpha\rangle$
-Proof see Appendix B
-We implemented our CHG Shapley data valuation method (see Algorithm 1) based on the code foundation of OpenDataVal (Jiang et al., 2023) and conducted experiments on the CIFAR-10 embedding dataset on a NVIDIA A40 server.
-The model used is pre-trained ResNet50 + logistic regression, and we report the results in Table 1.
-For noisy label detection tasks and comparisons with other data evaluation methods, see Picture 2. For point removal tasks, see Picture 1.
-The experimental results demonstrate that CHG Shapley data valuation method not only exhibits excellent data valuation and noise data detection capabilities, but also achieves remarkable operational efficiency.
-This is made possible by our research into the utility function that measures the impact of data subsets on model accuracy.
-In the next section, we will further show its data selection ability based on data valuation and noise label detection.
+Proof see Appendix B.
+The use of CHG Shapley to calculate the data value of each training data point is outlined in Algorithm 1, and the experimental results for different data valuation methods are presented in Table 1.
+Detailed experimental results comparing these data valuation methods can be found in Appendix A.
+Differences from Data Shapley: Although our approach shares the same underlying principle as Data Shapley, we do not try to approximate it.
+Data Shapley utilizes a model trained on a subset of the training set, evaluated on an additional validation set, as its utility function.
+In CHG Shapley, our utility function captures the extent of training loss reduction when using a subset of the training dataset.
+Furthermore, in Data Shapley, the required size of the validation set can grow by orders of magnitude to achieve the desired utility precision (i.e., the model’s accuracy on the validation set).
+For example, in a classification task, achieving a precision of 0.001 for the utility function may require 1,000 validation samples, while a precision of 0.0001 could demand 10,000 samples.
+Another challenge lies in the noise affecting the utility function, which arises from the model training process—particularly due to stochastic gradient descent—as discussed in Data Banzhaf (Wang & Jia, 2023).
+In contrast, CHG Shapley relies solely on the training set and incorporates both loss and gradient information from the model during training.
+This richer and more precise information allows CHG Shapley to achieve lower computational costs and higher precision.
+Moreover, by averaging the Shapley values over multiple epochs, CHG Shapley is somewhat resistant to the impacts of utility noise.
 Algorithm 1 CHG Shapley-based Data Valuation Algorithm
-0:  Training dataset ($N$), model architecture ($\theta_{0}$)
-0:  Value of every training data
-1:  Initialization: Model parameter and training hyperparameters initialization
-2:  for $k$-th training epoch do
-3:     Acquire loss and last-layer gradients of each data, i.e., $l_{i},\nabla f(i;\theta_{k})$
-4:     Calculate Shapley value of every data point based on Equation 3, with $\alpha=\frac{1}{N}\sum_{i\in N}l_{i}\nabla f(i;\theta)$ and $x_{i}=l_{i}\nabla f(i;\theta_{k})$
-5:     Update model parameters $\theta_{k}$ to $\theta_{k+1}$
-6:  end for
-7:  Return each data’s average Shapley value
+0:  Training dataset $N$, initial model parameters $\theta_{0}$, total epochs $K$
+0:  Shapley value of each training data point in $N$
+1:  for $k$-th epoch do
+2:     Acquire the loss and gradients for each data point in $N$, i.e., $h_{i},\nabla f(i;\theta_{k})$
+3:     Using Equation 4 to calculate $\phi_{i}^{k}$, the Shapley value of $i$-th data point in $k$ epoch, where $\alpha=\frac{1}{N}\sum_{i\in N}h_{i}\nabla f(i;\theta)$ and $x_{i}=h_{i}\nabla f(i;\theta_{k})$
+4:     Train model on $N$, updating parameters to $\theta_{k+1}$
+5:  end for
+6:  Return the average Shapley value for each data point across epochs: $\sum_{k=1}^{K}\phi_{i}^{k}/K$
 4 Data Selection
-In this section, we introduce various implementation strategies and practical techniques to enhance the scalability and efficiency of CHG Shapley. The experiments are conducted on a NVIDIA A40 server.
-Last-Layer Gradients: Modern deep learning models possess a vast number of parameters, resulting in high-dimensional gradients.
-To mitigate this issue, we adopt a last-layer gradient approximation as in(Killamsetty et al., 2021a, b).
-This technique focuses solely on the gradients of the last layer for neural networks in CHG Shapley, significantly improving the speed of the algorithm and other baseline methods.
-Per-Class Approximations: To enhance model performance, we adopt a per-class approach.
-This method involves caluclate the CHG Shapley separately for each class, which means considering only the data instances belonging to each class in each iteration.
-This approach not only reduces memory requirements but also decreases computational overhead, thereby improving the overall efficiency of the CHG Shapley-based data selection algorithm.
-CHG Shapley-based Data Selection Algorithm: Here, we detail the CHG Shapley Data Selection Algorithm, integrating the aforementioned techniques to ensure efficient and scalable data valuation and selection.
+4.1 From Data Valuation to Data Selection
+The Shapley value is a widely used method for quantifying individual contributions to overall utility in cooperative games.
+Previous work (Ghorbani & Zou, 2019; Wang & Jia, 2023) has also heuristically prioritized data points with high Shapley values for model training.
+In this section, we aim to lay the theoretical groundwork for employing Shapley values in data selection.
+Theorem 2
+Let $N$ be a set of players and $U$ a utility function.
+The Shapley value of player $i\in N$ is denoted by $\phi_{i}(U)$.
+Suppose that for all $i$ and for all subsets $S\subseteq N\setminus\{i\}$, the utility difference satisfies $-m\leq U(S\cup\{i\})-U(S)\leq M$.
+Then:
+  1. 1.
+    If $\phi_{i}(U)\geq 0$, then
+    $P_{\pi\sim\Pi}\left[U(S_{\pi}^{i}\cup\{i\})\geq U(S_{\pi}^{i})\right]\geq\frac{\phi_{i}(U)}{M},$
+  2. 2.
+    If $\phi_{i}(U)\leq 0$, then
+    $P_{\pi\sim\Pi}\left[U(S_{\pi}^{i}\cup\{i\})\geq U(S_{\pi}^{i})\right]\leq 1+\frac{\phi_{i}(U)}{m}.$
+Here, $S_{\pi}^{i}$ represents the set of players preceding player $i$ in a uniform random permutation $\pi$.
+The proof can be found in Appendix C.
+The theorem offers a probabilistic interpretation of Shapley values.
+Supposing $S_{\pi}^{i}$ represents the dataset selected so far to some extent, then, if $i$ has a positive Shapley value, the probability that it positively contributes to the coalition $S_{\pi}^{i}$ has a lower bound that increases proportionally with its Shapley value.
+Conversely, if $i$ has a negative Shapley value, the upper bound on this probability decreases as its Shapley value becomes smaller.
+Thus, when $i$ has a larger Shapley value, adding it to the selected training subset $S_{\pi}^{i}$ is more likely to yield a positive benefit for $S_{\pi}^{i}$.
+And when $S_{\pi}^{i}$ can not fairly approximate the dataset selected so far, using the original Shapley value for data selection may become ineffective.
 Algorithm 2 CHG Shapley-based Data Selection Algorithm
-0:  Training dataset ($N$), model architecture ($\theta_{0}$), selection interval $R=20$, the fraction (denoted as $a$) of data available for each epoch.
-0:  Model parameters
-1:  Initialization: Model parameter and training hyperparameters initialization, and let the training subset $S=\emptyset$.
-2:  for $k$-th training epoch do
+0:  Training dataset $N$, initial parameters $\theta_{0}$, total epochs $K$, selection interval $R=20$, fraction $a$ of selected data.
+0:  Final parameters $\theta_{K}$
+1:  Initialize subset $S=\emptyset$.
+2:  for $k$-th epoch do
 3:     if  $k\mod R==0$ then
-4:        Let training subset $S=\emptyset$
+4:        $S=\emptyset$
 5:        for $c$-th class do
-6:           Acquire loss and last-layer gradients of each data in this class, i.e., $l_{i},\nabla f(i;\theta_{k})$
-7:           Calculate Shapley value of every data point in this class (denoted as $N_{c}$) based on Equation 3, with $\alpha=\frac{1}{N_{c}}\sum_{i\in N_{c}}l_{i}\nabla f(i;\theta)$ and $x_{i}=l_{i}\nabla f(i;\theta_{k})$
-8:           Add the top $aN_{c}$ data points with the highest Shapley values to the training subset $S$
+6:           Acquire loss and last-layer gradients for $N_{c}$ (the subset of $N$ with label $c$), i.e., $h_{i},\nabla f(i;\theta_{k})$
+7:           Using Equation 4 to calculate Shapley values with $\alpha=\frac{1}{N_{c}}\sum_{i\in N_{c}}h_{i}\nabla f(i;\theta)$ and $x_{i}=h_{i}\nabla f(i;\theta_{k})$
+8:           Add top $aN_{c}$ points with highest Shapley values to $S$
 9:        end for
-10:        Calculate the Shapley values of each data point in $S$ with $\alpha=\frac{1}{S}\sum_{i\in S}l_{i}\nabla f(i;\theta)$ and $x_{i}=l_{i}\nabla f(i;\theta_{k})$, and apply minmax normalization to Shapley values to get the weights (denoted as $\{w_{S}\}$) of each data point in $S$
-11:     end if
-12:     Train model on $S$ using weighted learning with $\{w_{S}\}$, updating its parameters to $\theta_{k+1}$
-13:  end for
-Results: We report the accuracy and time consumption comparisons in Table 2 under the data pruning setting.
-CHG Shapley demonstrates particularly strong performance when the selection ratio is small, highlighting its ability to identify high-value data effectively.
-Additionally, the time consumption remains acceptable even when considering data selection processes, underscoring the high efficiency of CHG Shapley as a data valuation method.
-This efficiency, combined with its accuracy, verifies that CHG Shapley is a practical and powerful tool for data selection and valuation in large-scale machine learning tasks.
-The results presented in Table 3, where 30% of the class labels were randomly replaced with other labels, provide additional evidence of CHG Shapley’s capability to detect mislabeled data.
-CHG Shapley demonstrates superior performance compared to other methods in this scenario, reinforcing its effectiveness in identifying and addressing label noise.
-Table 2: The accuracy (%) and time (s) comparison. All methods are trained with ResNet-18.
-Dataset CIFAR10 CIFAR100 Tiny ImageNet Fraction 0.05 0.1 0.3 0.5 0.7 0.05 0.1 0.3 0.5 0.7 0.05 0.1 0.3 0.5 0.7 Accuracy (%) Full 95.51 77.56 62.43 Random 65.90 77.36 89.93 93.11 93.87 22.58 36.51 62.27 69.70 73.04 17.13 27.70 45.17 52.37 57.89 AdaptiveRandom 83.26 88.78 93.88 94.77 94.66 48.52 60.58 72.45 74.97 75.70 37.76 48.50 58.49 60.30 62.18 Glister 85.44 89.92 91.26 91.93 93.04 48.25 64.63 72.72 74.00 75.14 37.17 48.29 56.77 59.13 61.50 GradMatch (PB) 85.17 90.17 93.93 94.42 94.83 50.42 63.53 73.79 75.78 76.20 38.88 49.97 59.79 61.56 62.19 CHG Shapley 87.15 90.88 93.65 94.51 94.58 53.97 64.46 72.54 74.38 75.71 38.92 49.00 58.44 60.68 61.80 Hardness Shapley 54.96 68.21 84.18 87.02 - 34.54 53.11 66.20 69.10 - - - - - - Gradient Shapley 84.87 89.95 93.54 94.08 - 51.15 59.73 70.69 73.88 - - - - - - Time (s) Full 7745 8082 49747 Random 776 1140 1906 6489 8708 805 1645 3839 3170 5859 2742 5184 15323 24475 33910 AdaptiveRandom 502 1522 4037 6815 6574 746 1358 3324 4786 7166 3451 5852 21366 27659 37583 Glister 1600 2736 4713 11853 15740 1247 1885 4778 11762 14611 4382 7693 19210 28289 36558 GradMatch (PB) 1308 2129 4897 7661 9916 804 2192 4077 6095 9185 4455 6904 24238 30814 51780 CHG Shapley 1383 1768 3746 7356 9526 1258 2037 4706 7012 9657 7077 9887 25682 30762 38177 Hardness Shapley 1191 1741 3606 4878 - 1233 1836 3828 5824 - - - - - - Gradient Shapley 1365 1933 2714 4017 - 1427 2025 2730 4032 - - - - - -
-Table 3: Comparison of performance under noise label setting.
-Dataset CIFAR10 CIFAR100 Fraction 0.05 0.1 0.3 0.5 0.7 0.05 0.1 0.3 0.5 0.7 AdaptiveRandom 57.18 63.21 71.84 75.38 77.41 24.12 32.69 45.36 52.18 53.00 Glister 61.17 70.39 73.73 75.64 76.59 18.64 29.22 45.80 49.78 52.45 GradMatch (PB) 58.38 69.67 72.59 75.41 76.90 25.90 33.84 47.71 52.35 54.29 CHG Shapley 67.43 77.32 85.33 88.83 88.90 35.75 46.41 59.59 61.92 63.36
+10:     end if
+11:     Train model on $S$, updating parameters to $\theta_{k+1}$
+12:  end for
+4.2 CHG Shapley-based Data Selection Algorithm
+In this section, we introduce several implementation strategies and practical techniques to improve the scalability and efficiency of CHG Shapley, incorporating common tricks from data selection methods (Killamsetty et al., 2021a; b), and all the compared methods (except Full, Random, and AdaptiveRandom) also adopt these tricks:
+Interval Data Selection: Instead of selecting new data at every training step, we perform data selection at intervals.
+For example, during a 300-epoch training process, data selection is done every 20 epochs, while in the remaining epochs, the previously selected subset is used for training.
+This approach significantly reduces the computational overhead associated with data selection.
+Last-Layer Gradients: Given the high dimensionality of gradients in modern deep learning models, we adopt a last-layer gradient approximation.
+By focusing solely on the gradients of the last layer, this technique accelerates the computation for all methods.
+Per-Class Approximations: To further enhance model accuracy, we apply a per-class approach, running the algorithm separately for each class by considering only the data points from that class in each data selection iteration.
+This reduces both memory usage and computational cost, improving the overall performance of all methods.
+We detail the CHG Shapley Data Selection Algorithm in Algorithm 2, integrating the aforementioned techniques to ensure efficient and scalable data valuation-based data selection.
+5 Experiments
+We first present the data valuation experiments, followed by the data selection experiments.
+All experiments were conducted on a server equipped with an Intel(R) Xeon(R) Gold 6326 CPU (2.90 GHz) and an NVIDIA A40 GPU.
+All methods were implemented using PyTorch.
+5.1 Data Valuation Setup and Results
+Our CHG Shapley data valuation method (Algorithm 1) was implemented using the OpenDataVal framework (Jiang et al., 2023), and experiments were conducted on the CIFAR-10 embedding dataset.
+The model used for evaluation was a pre-trained ResNet50 followed by logistic regression, with results reported in Table 1.
+The methods compared include Leave-One-Out, Influence Subsample (an efficient approximation of the Influence function (Koh & Liang, 2017)), DVRL (Yoon et al., 2020), Data Banzhaf (Wang & Jia, 2023), AME (Lin et al., 2022), LAVA (Just et al., 2023), and Data OOB (Kwon & Zou, 2023).
+We do not include Data Shapley and Bata Shapley because these two methods are too slow.
+We evaluated performance on three tasks: noisy label detection (Table 5), noisy feature detection (Fig.
+3), and point removal experiments (Fig.
+3).
+Detailed results comparing these methods can be found in Appendix A.
+Across the three tasks, Data OOB and CHG Shapley ranked first and second, respectively.
+Notably, CHG Shapley completed its evaluation in 18 seconds, whereas Data OOB took over 2 hours.
+These results demonstrate that CHG Shapley not only achieves competitive data valuation and noise detection performance but also offers significant operational efficiency.
+This efficiency is driven by the analytic form of the proposed utility function, which evaluates the influence of data subsets on training loss.
+In the next section, we explore CHG Shapley’s effectiveness in data selection for larger, modern datasets.
+Table 2: The accuracy (%) and time (h) comparison. All methods are trained with ResNet-18.
+Dataset CIFAR10 CIFAR100 Fraction 0.05 0.1 0.3 0.5 0.7 0.05 0.1 0.3 0.5 0.7 Accuracy (%) Full 95.51 77.56 Random 66.63 77.36 89.93 93.11 93.87 22.58 36.51 62.27 69.70 73.04 AdaptiveRandom 83.26 88.78 93.88 94.58 94.66 48.52 60.58 72.45 74.97 75.70 Glister 85.42 90.52 89.80 90.15 93.36 46.54 63.88 71.88 73.93 75.25 GradMatchPB 84.91 90.14 93.73 94.44 94.54 49.01 62.56 71.27 73.66 74.43 Hardness Shapley 63.07 79.96 93.93 94.63 94.78 37.78 55.21 72.46 75.86 75.77 TracIn (Gradient-Dot) 63.47 81.49 94.13 94.85 94.62 40.21 56.55 72.11 75.01 76.33 CGSV (Gradient-Cosine) 72.97 83.49 92.24 93.68 93.98 44.08 53.87 68.16 74.04 75.59 GradE Shapley (This paper) 83.10 89.80 93.10 93.81 93.58 50.38 60.16 70.53 73.66 74.57 CHG Shapley (This paper) 85.47 91.20 93.73 94.42 94.07 53.20 63.70 72.32 74.59 75.31 Time (h) Full 2.2 2.2 Random 0.1 0.3 0.5 1.8 2.4 0.2 0.5 1.1 0.9 1.6 AdaptiveRandom 0.1 0.4 1.1 1.5 1.8 0.2 0.4 0.9 1.3 2.0 Glister 0.4 0.8 2.2 2.4 2.3 0.4 1.0 3.2 3.2 3.6 GradMatchPB 0.2 0.4 1.0 1.4 2.3 0.3 0.5 1.2 2.2 3.0 Hardness Shapley 0.3 0.6 1.0 1.8 2.4 0.4 0.8 1.3 2.9 3.6 TracIn (Gradient-Dot) 0.3 0.5 1.2 1.8 2.4 0.4 0.6 1.3 2.9 2.4 CGSV (Gradient-Cosine) 0.4 0.5 1.1 1.7 2.4 0.4 0.8 1.2 2.8 2.4 GradE Shapley (This paper) 0.3 0.6 1.2 1.8 2.7 0.4 0.5 1.3 1.9 3.6 CHG Shapley (This paper) 0.3 0.6 1.4 1.7 2.2 0.4 0.5 1.3 2.1 3.5
+5.2 Data selection setup
+In this experiment, we employed a ResNet-18 model, trained from scratch using stochastic gradient descent with a learning rate of 0.05, momentum of 0.9, weight decay set to 5e-4, and Nesterov acceleration.
+The model was trained over 300 epochs, with data reselection occurring every 20 epochs.
+The fraction of selected data varied across 0.05, 0.1, 0.3, 0.5, and 0.7.
+In the class imbalance experiments, we set the imbalance ratio to 0.3, meaning that for 30% of the classes, only 10% of their training and validation data was retained.
+In the noisy label experiments, we set the noise ratio to 0.3, meaning that 30% of the labels in the training set were randomly replaced with labels from the available classes.
+The compared methods include Full (training the model with the full dataset), Random (selecting data in the first epoch and never updating the selection), AdaptiveRandom (selecting data randomly in every selection epoch), Glister (Killamsetty et al., 2021b) and GradMatchPB (Killamsetty et al., 2021a) (two coreset selection methods), Hardness Shapley (selecting data with the highest loss), TracIn (Pruthi et al., 2020), and CGSV (Xu et al., 2021) (which use gradient inner product or cosine similarity for data selection as in (Xia et al., 2024)), as well as our proposed GradE Shapley and CHG Shapley.
+To ensure a fair comparison, the target gradient for TracIn, CGSV, GradE Shapley, and CHG Shapley is the mean gradient of all training data, eliminating the need for an additional validation set.
+All methods were implemented using the Cords framework (Killamsetty et al., 2021b; a).
+Table 3: Comparison of performance under class imbalance setting.
+Dataset CIFAR10 CIFAR100 Fraction 0.05 0.1 0.3 0.5 0.7 0.05 0.1 0.3 0.5 0.7 Full 90.37 65.11 AdaptiveRandom 73.64 82.00 88.73 89.61 89.92 36.45 49.44 61.38 63.44 64.27 Glister 67.80 81.47 83.16 82.07 88.63 32.79 48.15 60.86 61.94 62.72 GradMatchPB 75.12 82.92 88.21 89.55 89.71 39.45 51.39 61.54 63.97 64.07 Hardness Shapley 35.57 64.95 87.95 89.13 89.56 25.87 43.34 61.42 63.43 64.31 TracIn (Gradient-Dot) 42.97 68.34 88.45 89.42 89.54 31.81 43.52 61.10 63.54 64.02 CGSV (Gradient-Cosine) 63.09 74.45 87.60 89.72 89.01 32.35 42.58 58.03 63.0 64.38 GradE Shapley (This paper) 67.61 80.98 86.46 87.55 88.55 39.65 48.96 57.81 61.02 62.14 CHG Shapley (This paper) 69.94 83.65 87.69 88.61 88.86 40.31 51.88 60.33 62.90 63.89
+Table 4: Comparison of performance under noisy label setting.
+Dataset CIFAR10 CIFAR100 Fraction 0.05 0.1 0.3 0.5 0.7 0.05 0.1 0.3 0.5 0.7 Full 79.95 56.67 AdaptiveRandom 57.18 63.21 71.84 75.38 77.41 24.12 32.69 45.36 52.18 53.00 Glister 62.48 71.22 75.51 77.62 77.91 18.75 27.23 46.22 49.38 52.93 GradMatchPB 56.33 69.29 73.50 76.44 76.74 26.40 32.55 43.75 48.47 50.50 Hardness Shapley 21.63 35.41 63.31 75.95 77.28 3.72 7.23 43.16 50.90 51.80 TracIn (Gradient-Dot) 30.72 46.87 70.45 77.14 77.90 28.37 29.70 43.05 50.83 52.29 CGSV (Gradient-Cosine) 70.85 77.48 78.90 78.42 76.84 33.12 42.12 55.31 59.23 57.18 CHG Shapley (This paper) 72.18 83.35 88.57 90.00 90.20 36.75 46.93 60.09 62.59 62.79 GradE Shapley (This paper) 79.35 86.74 90.83 92.03 91.35 44.04 52.14 60.25 63.20 62.92
+5.3 Data Selection Results
+We present the accuracy and time consumption comparisons in Table 2 under the standard dataset setting.
+Hardness Shapley also performs competitively at higher fractions (e.g., 0.3), aligning with the findings of Qin et al.
+(2024).
+However, when the selection ratio is small (e.g., 0.05), its performance significantly declines, likely due to the difficulty in training the network when all selected data points are particularly hard.
+TracIn also performs well when the selection ratio is large.
+CHG Shapley demonstrates particularly strong performance when the selection ratio is small (i.e., 0.05, 0.1), effectively identifying high-value data.
+Furthermore, its time consumption is comparable to other methods, underscoring the efficiency of CHG Shapley as a data valuation approach.
+Table 3 presents the results under the class imbalance dataset setting.
+Similar to the standard dataset setting, CHG Shapley performs well at smaller selection ratios, and its comparison with GradE Shapley highlights the advantage of incorporating data hardness as part of data value.
+The most surprising results are presented in Table 4, under the noisy label dataset setting.
+GradE Shapley demonstrates a 5%-10% performance improvement over other methods (except CHG Shapley) in this scenario.
+Even with 10%-30% of the data selected, GradE Shapley can outperform using the full dataset, reinforcing its effectiveness in identifying label noise.
+In contrast, Hardness Shapley performs notably worse in this setting, especially at smaller selection ratios.
+In other words, the unreliability of hardness in the presence of label noise leads to CHG Shapley underperforming relative to GradE Shapley.
+Overall, the combination of efficiency and accuracy demonstrated across these three experiments establishes CHG Shapley as a practical and powerful tool for data selection and valuation in large-scale datasets with small amounts of label noise.
+And GradE Shapley excels when handling larger amounts of label noise.
 6 Discussion and Future Work
-In this paper, we define the Conduct of Hardness and Gradient score (CHG score) of a data subset $S$ as $\|\frac{1}{N}\sum_{i\in N}l_{i}\nabla f(i;\theta)\|_{2}^{2}-\|\frac{1}{N}\sum_{i\in N}l_{i}\nabla f(i;\theta)-\frac{1}{S}\sum_{i\in S}l_{i}\nabla f(i;\theta)\|_{2}^{2}$, and use this score based on Euclidean distance to measure the influence of data subsets on model accuracy.
-However, it’s worth noting that many studies on data selection based on gradients utilize different distance measures between $\frac{1}{N}\sum_{i\in N}l_{i}\nabla f(i;\theta)$ and $\frac{1}{S}\sum_{i\in S}l_{i}\nabla f(i;\theta)$, such as cosine similarity (Zhao et al., 2021), inner product, L1 norm, or their combinations, and then optimize this metric to select data.
-However, whether these metrics can be used for data valuation remains unknown because, unlike the theoretical support for the aforementioned Euclidean distance, these metrics lack theoretical grounding.
-Whether cosine similarity, inner product, L1 norm, or their combinations can achieve this goal remains to be investigated.
-On the other hand, if cosine similarity and L1 norm are used in our framework, it becomes challenging to derive the corresponding Shapley values.
-In other words, if the Euclidean distance cannot effectively measure the utility of dataset subsets for model accuracy, the entire data valuation approach proposed in this study becomes computationally infeasible.
-The root of this problem lies in the complexity of computing Shapley values, which is generally challenging and only tractable when the utility function follows a specific form (such as linear models, Euclidean distance, or inner product).
+In the data selection experiments, the introduction of hardness improves the model’s performance when no label noise is present.
+This suggests that hardness can serve as a proxy for data value to some extent.
+However, in cases with significant label noise, incorporating hardness into GradE Shapley diminishes its effectiveness.
+This indicates a need for further investigation into how hardness can be fairly integrated into data value assessments, particularly when it becomes an unreliable metric.
+Besides, CHG Shapley performs well at small fractions but struggles with larger ones.
+This challenge may arise from the difficulties of selecting data based on Shapley values (Wang et al., 2024c), rather than the limitations of CHG Shapley itself.
+Specifically, once a subset of data $s$ is selected, the Shapley value of the remaining data may change, as the set $S_{\pi}^{i}$ in Equation 2 will include $s$.
+This affects the effectiveness of the original Shapley values in evaluating the remaining data.
+If we must recalculate the Shapley values for the remaining data each time a new data point is selected, the time complexity becomes unmanage able even with our proposed method.
+Thus, we leave the exploration of iterative Shapley value calculations for data selection to future work.
 Appendix A Data Valuation results
+We adhere to the experimental settings and implementation described in (Jiang et al., 2023) to conduct the data valuation experiment.
+We evaluate several data valuation methods on the CIFAR10-embeddings dataset, using a pretrained ResNet50 model for feature extraction, followed by a logistic regression model for classification.
+The dataset contains 1,000 training samples, 100 validation samples, and 300 test samples.
+The noise rate is set as 0.1, which means 10% of the labels in both the training and validation datasets will be randomly altered.
+We assess the performance using classification accuracy.
+The logistic regression model is trained for 10 epochs with a batch size of 100 and a learning rate of 0.01.
+The data valuation methods compared in this study include CHG Shapley, Leave-One-Out, Influence Subsample (an efficient approximation of the Influence function), DVRL, Data Banzhaf, AME, LAVA, and Data OOB.
+We have excluded the results for Data Shapley and Beta Shapley, as these two methods are at least 10 times more time-consuming than the methods evaluated.
+The runtime comparison results are provided in Table 6.
+A.1 Noisy label detection
+We conduct noisy data detection by introducing mislabeled data into the dataset and applying various data valuation algorithms to identify the mislabeled instances.
+The performance of each algorithm is measured using the F1-score, which evaluates the balance between precision and recall.
+A higher F1-score indicates better accuracy in detecting mislabeled data.
+As shown in Table 5, Data OOB achieves the highest F1-score, while CHG Shapley follows closely as the second-best performer.
+This indicates that CHG Shapley demonstrates competitive noisy data detection capabilities compared to state-of-the-art data valuation methods.
+Table 5: F1-scores of various data valuation methods for noisy data detection
+Method F1-Score
+Leave-One-Out 0.184573
+Random Evaluator 0.173623
+AME (1000 models) 0.181989
+DVRL (2000 rl epochs) 0.297872
+Data Banzhaf (1000 models) 0.160883
+Data OOB (1000 models) 0.368821
+Influence Subsample (1000 models) 0.199662
+Lava Evaluator 0.212329
+CHG Shapley 0.334405
+A.2 Noisy Feature detection
+In Fig.
+3, we visualize the effectiveness of each data valuation method in identifying noisy data points by inspecting a fraction of the dataset.
+The x-axis represents the proportion of inspected data, while the y-axis indicates the proportion of identified corrupted data.
+The orange curve represents the optimal performance, while the blue curve reflects the true performance of the evaluator.
+A closer alignment of the blue curve with the orange curve indicates a more effective data valuation method.
+Notably, CHG Shapley and Data OOB emerge as the top performers, demonstrating their superior ability to detect noisy samples compared to other evaluated methods.
+Figure 2: Noisy feature detection experiment on CIFAR10-embeddings dataset.
 Figure 3: Point removal experiment on CIFAR10-embeddings dataset.
-Figure 2: Discover noisy datas experiment on CIFAR10-embeddings dataset.
-Appendix B Proof of Theorem  1
-From $U(S)=||\alpha||^{2}-||\frac{1}{S}\sum_{i\in S}x_{i}-\alpha||^{2}$ we have $U(S)=-||\frac{1}{S}\sum_{i\in S}x_{i}||^{2}+2\frac{1}{S}\sum_{i\in S}x_{i}\alpha=U_{1}(S)+U_{2}(S)$, where $U_{1}(S)=-||\frac{1}{S}\sum_{i\in S}x_{i}||^{2}$, $U_{2}(S)=2\frac{1}{S}\sum_{i\in S}x_{i}\alpha$.
-For $U_{2}$, we have: $\phi_{j}(U_{2})=\left(\frac{1}{n}\sum_{k=1}^{n}\frac{1}{k}\right)x_{j}\alpha-\left(\frac{1}{n}\sum_{k=2}^{n}\frac{1}{k}\right)(\frac{1}{n-1}\sum_{i\in N/{j}}x_{i}\alpha)$
-Then we have: $\phi_{j}(U_{2})=\left(\frac{1}{n-1}\sum_{k=1}^{n-1}\frac{1}{k}\right)x_{j}\alpha-\left(\frac{1}{n-1}\sum_{k=2}^{n}\frac{1}{k}\right)(\frac{1}{n}\sum_{i\in N}x_{i}\alpha)$
-Then:
-$U_{1}(S\cup j)-U_{1}(S)=-\frac{1}{(S+1)^{2}}x_{j}^{2}-\frac{2}{(S+1)^{2}}x_{j}(\sum_{i\in S}x_{i})+\frac{2S+1}{S^{2}(S+1)^{2}}(\sum_{i\in S}x_{i})^{2}.$
-Because of the linearity of the Shapley Value, we can calculate the Shapley value under utility functions $U_{1}$ and $U_{2}$ separately, and $\phi_{j}(U)=\phi_{j}(U_{1})+\phi_{j}(U_{2})$.
+A.3 Point Removal Experiment
+In this experiment, we evaluate the impact of removing data points based on the rankings generated by each data valuation algorithm.
+After each removal of high- or low-value data points, we retrain the model and assess its performance on a test dataset.
+The x-axis represents the fraction of removed high or low-value data, while the y-axis indicates the accuracy of the retrained model using the remaining data.
+Two curves are generated: one illustrating performance when data is removed in descending order (orange) and the other in ascending order (blue).
+For the orange curve, a lower accuracy indicates better ability to detect high-value data, while a higher accuracy is preferable for the blue curve cause removing low-value data should have little impact to model accuracy.
+As shown in Fig.
+3, both CHG Shapley and Data OOB perform exceptionally well in this experiment, reinforcing their positions as leading methods for effective data valuation and highlighting their capacity to optimize the removal of noisy samples for enhanced model performance.
+A.4 runtime comparison
+The runtime comparison in Table 6 shows that CHG Shapley is one of the fastest methods, completing in just 18 seconds, while methods like Leave-One-Out and Data OOB take over an hour.
+Data Shapley and Beta Shapley, on the other hand, are far slower, requiring more than 24 hours.
+This highlights CHG Shapley’s efficiency for practical use.
+Table 6: Comparison of runtimes for different data valuation methods on CIFAR10-embeddings
+Method Elapsed Time (hh:mm:ss)
+Leave-One-Out 1:41:25
+Influence Subsample (1000 models) 1:11:42
+DVRL(2000 rl epochs) 00:06:10
+Data Banzhaf (1000 models) 1:04:06
+AME (1000 models) 4:38:19
+Data OOB (1000 models) 2:02:03
+Data Shapley (1000 models) More than 24h
+Beta Shapley (1000 models) More than 24h
+Lava Evaluator 00:00:02
+CHG Shapley 00:00:18
+Appendix B Proof of Theorem 1
+Because $U(S)=||\alpha||^{2}-||\frac{1}{|S|}\sum_{i\in S}x_{i}-\alpha||^{2}=-||\frac{1}{|S|}\sum_{i\in S}x_{i}||^{2}+2\frac{1}{|S|}\sum_{i\in S}\langle x_{i},\alpha\rangle$, and we define $U_{1}(S)=-||\frac{1}{|S|}\sum_{i\in S}x_{i}||^{2}$ and $U_{2}(S)=2\frac{1}{|S|}\sum_{i\in S}x_{i}\alpha$, then $U(S)=U_{1}(S)+2U_{2}(S)$.
+The basic idea here is to use the Linearity property of the Shapley value.
+By calculating $\phi_{j}(U_{1})$ and $\phi_{j}(U_{2})$ separately, we can get $j$-th player’s Shapley value as: $\phi_{j}(U)=\phi_{j}(U_{1})+2\phi_{j}(U_{2})$.
+For $U_{2}$, we have:
+$U_{2}(S\cup j)-U_{2}(S)$ $=\frac{1}{|S|+1}\sum_{i\in S\cup j}\langle x_{i},\alpha\rangle-\frac{1}{|S|}\sum_{i\in S}\langle x_{i},\alpha\rangle$
+$=\frac{\langle x_{j},\alpha\rangle}{|S|+1}-\frac{\sum_{i\in S}\langle x_{i},\alpha\rangle}{|S|(|S|+1)}.$
+Then we have:
+$\phi_{j}(U_{2})$ $=\frac{1}{n}\sum_{k=1}^{n}{n-1\choose k-1}^{-1}\sum_{S\subseteq N\setminus\{j\},|S|=k-1}\left[U_{2}(S\cup j)-U_{2}(S)\right]$
+$=\frac{1}{n}\sum_{k=1}^{n}{n-1\choose k-1}^{-1}\sum_{\begin{subarray}{c}S\subseteq N\setminus\{j\},\\ |S|=k-1\end{subarray}}\frac{\langle x_{j},\alpha\rangle}{|S|+1}-\frac{1}{n}\sum_{k=1}^{n}{n-1\choose k-1}^{-1}\sum_{\begin{subarray}{c}S\subseteq N\setminus\{j\},\\ |S|=k-1\end{subarray}}\frac{\sum_{i\in S}\langle x_{i},\alpha\rangle}{|S|(|S|+1)}$
+$=\frac{1}{n}\sum_{k=1}^{n}{n-1\choose k-1}^{-1}{n-1\choose k-1}\frac{\langle x_{j},\alpha\rangle}{k}-\frac{1}{n}\sum_{k=2}^{n}{n-1\choose k-1}^{-1}\sum_{i\in N\setminus\{j\}}\sum_{\begin{subarray}{c}S\subseteq N\setminus\{i,j\},\\ |S|=k-2\end{subarray}}\frac{\langle x_{i},\alpha\rangle}{(k-1)k}$
+$=\frac{1}{n}\sum_{k=1}^{n}\frac{\langle x_{j},\alpha\rangle}{k}-\frac{1}{n}\sum_{k=2}^{n}{n-1\choose k-1}^{-1}\sum_{i\in N\setminus\{j\}}{n-2\choose k-2}\frac{\langle x_{i},\alpha\rangle}{(k-1)k}$
+$=\frac{\sum_{k=1}^{n}\frac{1}{k}}{n}\langle x_{j},\alpha\rangle-\frac{1}{n}\sum_{k=2}^{n}\frac{k-1}{n-1}\sum_{i\in N\setminus\{j\}}\frac{\langle x_{i},\alpha\rangle}{(k-1)k}$
+$=\frac{\sum_{k=1}^{n}\frac{1}{k}}{n}\langle x_{j},\alpha\rangle-\frac{\sum_{k=2}^{n}\frac{1}{k}}{n(n-1)}\langle\sum_{i\in N\setminus\{j\}}x_{i},\alpha\rangle$
+$=\frac{\sum_{k=1}^{n}\frac{1}{k}-\frac{1}{n}}{n-1}\langle x_{j},\alpha\rangle-\frac{\sum_{k=1}^{n}\frac{1}{k}-1}{n(n-1)}\langle\sum_{i\in N}x_{i},\alpha\rangle.$
+Thus, $\phi_{j}(U_{2})=\frac{\sum_{k=1}^{n}\frac{1}{k}-\frac{1}{n}}{n-1}\langle x_{j},\alpha\rangle-\frac{\sum_{k=1}^{n}\frac{1}{k}-1}{n(n-1)}\langle\sum_{i\in N}x_{i},\alpha\rangle$.
+Similarily, for $U_{1}$, we have:
+$U_{1}(S\cup j)-U_{1}(S)=-\frac{1}{(|S|+1)^{2}}x_{j}^{2}-\frac{2}{(|S|+1)^{2}}x_{j}(\sum_{i\in S}x_{i})+\frac{2|S|+1}{|S|^{2}(|S|+1)^{2}}(\sum_{i\in S}x_{i})^{2}.$
+Then we have:
 $\phi_{j}(U_{1})$ $=-\left(\frac{1}{n}\sum_{k=1}^{n}\frac{1}{k^{2}}\right)x_{j}^{2}-\left(\frac{2}{n(n-1)}\sum_{k=2}^{n}\frac{(k-1)}{k^{2}}\right)(\sum_{i\in N/{j}}x_{i})x_{j}$
@@
 $+\left(\frac{1}{n(n-1)(n-2)}\sum_{k=3}^{n}\frac{(2k-1)(k-2)}{k^{2}(k-1)}\right)(\sum_{a\in N/{j}}\sum_{b\in N/{j},a\neq b}x_{a}x_{b})$
-Then we have:
+Then:
 $\phi_{j}(U_{1})$ $=-\left(\frac{1}{n}\sum_{k=1}^{n}\frac{1}{k^{2}}\right)x_{j}^{2}+\left(\frac{2}{n(n-1)}\sum_{k=2}^{n}\frac{(k-1)}{k^{2}}\right)x_{j}^{2}-\left(\frac{2}{n(n-1)}\sum_{k=2}^{n}\frac{(k-1)}{k^{2}}\right)(\sum_{i\in N}x_{i})x_{j}$
@@
 $+\left(\frac{1}{n(n-1)(n-2)}\sum_{k=3}^{n}\frac{(2k-1)(k-2)}{k^{2}(k-1)}\right)(\sum_{a\in N}\sum_{b\in N}x_{a}x_{b})$
-$+\left(\frac{1}{n(n-1)}\sum_{k=2}^{n}\frac{(2k-1)}{k^{2}(k-1)}-\frac{1}{n(n-1)(n-2)}\sum_{k=3}^{n}\frac{(2k-1)(k-2)}{k^{2}(k-1)}\right)(\sum_{i\in N}x_{i}^{2})$
-Then we have:
+$+\left(\frac{1}{n(n-1)}\sum_{k=2}^{n}\frac{(2k-1)}{k^{2}(k-1)}-\frac{1}{n(n-1)(n-2)}\sum_{k=3}^{n}\frac{(2k-1)(k-2)}{k^{2}(k-1)}\right)(\sum_{i\in N}x_{i}^{2}),$
+And:
 $\phi_{j}(U_{1})$
@@
 $+\frac{1}{n(n-1)(n-2)}\left(2\sum_{k=2}^{n}\frac{1}{k}-2\sum_{k=2}^{n}\frac{1}{k^{2}}-1+\frac{1}{n}\right)(\sum_{a\in N}\sum_{b\in N}x_{a}x_{b})$
-$+\left(\frac{1}{n(n-1)}\left(\sum_{k=2}^{n}\frac{1}{k^{2}}+1-\frac{1}{n}\right)-\frac{2\sum_{k=2}^{n}\frac{1}{k}-2\sum_{k=2}^{n}\frac{1}{k^{2}}-1+\frac{1}{n}}{n(n-1)(n-2)}\right)(\sum_{i\in N}x_{i}^{2})$
-Then we have:
+$+\left(\frac{1}{n(n-1)}\left(\sum_{k=2}^{n}\frac{1}{k^{2}}+1-\frac{1}{n}\right)-\frac{2\sum_{k=2}^{n}\frac{1}{k}-2\sum_{k=2}^{n}\frac{1}{k^{2}}-1+\frac{1}{n}}{n(n-1)(n-2)}\right)(\sum_{i\in N}x_{i}^{2}),$
+Thus:
 $\phi_{j}(U_{1})$
@@
 $+\frac{1}{n(n-1)(n-2)}\left(2\sum_{k=1}^{n}\frac{1}{k}-2\sum_{k=1}^{n}\frac{1}{k^{2}}-1+\frac{1}{n}\right)(\sum_{a\in N}\sum_{b\in N}x_{a}x_{b})$
-$+\left(\frac{1}{n(n-1)}\left(\sum_{k=1}^{n}\frac{1}{k^{2}}-\frac{1}{n}\right)-\frac{2\sum_{k=2}^{n}\frac{1}{k}-2\sum_{k=2}^{n}\frac{1}{k^{2}}-1+\frac{1}{n}}{n(n-1)(n-2)}\right)(\sum_{i\in N}x_{i}^{2})$
-so :
+$+\left(\frac{1}{n(n-1)}\left(\sum_{k=1}^{n}\frac{1}{k^{2}}-\frac{1}{n}\right)-\frac{2\sum_{k=2}^{n}\frac{1}{k}-2\sum_{k=2}^{n}\frac{1}{k^{2}}-1+\frac{1}{n}}{n(n-1)(n-2)}\right)(\sum_{i\in N}x_{i}^{2}),$
+Finally, we get the $j$-th player’s Shapley value:
 $\phi_{j}(U)$
@@
 $+\frac{2}{n-1}\left(\sum_{k=1}^{n}\frac{1}{k}-\frac{1}{n}\right)x_{j}\alpha-\frac{2}{n(n-1)}\left(\sum_{k=1}^{n}\frac{1}{k}-1\right)(\sum_{i\in N}x_{i})\alpha$
+Appendix C Proof of Theorem  2
+Theorem 2: Let $N$ be a set of players and $U$ a utility function.
+The Shapley value of player $i\in N$ is denoted by $\phi_{i}(U)$.
+Suppose that for all $i$ and for all subsets $S\subseteq N\setminus\{i\}$, the utility difference satisfies $-m\leq U(S\cup\{i\})-U(S)\leq M$.
+Then:
+  1. 1.
+    If $\phi_{i}(U)\geq 0$, then
+    $P_{\pi\sim\Pi}\left[U(S_{\pi}^{i}\cup\{i\})\geq U(S_{\pi}^{i})\right]\geq\frac{\phi_{i}(U)}{M},$
+  2. 2.
+    If $\phi_{i}(U)\leq 0$, then
+    $P_{\pi\sim\Pi}\left[U(S_{\pi}^{i}\cup\{i\})\geq U(S_{\pi}^{i})\right]\leq 1+\frac{\phi_{i}(U)}{m}.$
+Here, $S_{\pi}^{i}$ represents the set of players preceding player $i$ in a uniform random permutation $\pi$.
+Our proof strategy is first establishing a variant of Markov’s inequality, then substituting $U(S_{\pi}^{i}\cup\{i\})-U(S_{\pi}^{i})$ as the variable into the inequality.
+C.1 A variant of Markov’s inequality
+  1. 1.
+    For $\mathbb{E}[X]\geq 0$:
+    Since $X\cdot I[X<0]\leq 0$, we have $\mathbb{E}[X\cdot I[X<0]]\leq 0$. Therefore, we can state:
+    $\mathbb{E}[X]\leq\mathbb{E}[X\cdot I[X\geq 0]].$
+    Given that $X\leq M$, we obtain:
+    $X\cdot I[X\geq 0]\leq M\cdot I[X\geq 0].$
+    Thus,
+    $\mathbb{E}[X\cdot I[X\geq 0]]\leq\mathbb{E}[M\cdot I[X\geq 0]]=M\cdot P[X\geq 0].$
+    This leads us to:
+    $\mathbb{E}[X]\leq M\cdot P[X\geq 0].$
+    Rearranging gives:
+    $P[X\geq 0]\geq\frac{\mathbb{E}[X]}{M}.$
+  2. 2.
+    For $\mathbb{E}[X]\leq 0$:
+    Similarly, since $X\cdot I[X\geq 0]\geq 0$, it follows that $\mathbb{E}[X\cdot I[X\geq 0]]\geq 0$. Therefore, we have:
+    $\mathbb{E}[X]\geq\mathbb{E}[X\cdot I[X<0]].$
+    Since $X\geq-m$, we find:
+    $X\cdot I[X<0]\geq-m\cdot I[X<0].$
+    Thus,
+    $\mathbb{E}[X\cdot I[X<0]]\geq\mathbb{E}[-m\cdot I[X<0]]=-m\cdot P[X<0].$
+    This allows us to write:
+    $\mathbb{E}[X]\geq-m\cdot P[X<0].$
+    Rearranging leads to:
+    $P[X<0]\leq\frac{\mathbb{E}[X]}{-m}.$
+    Consequently, we have:
+    $P[X\geq 0]\geq 1+\frac{\mathbb{E}[X]}{m}.$
+C.2 Proof
+Let $X:=U(S_{\pi}^{i}\cup\{i\})-U(S_{\pi}^{i})$, then the expectation of $X$ over the uniformly random permutation $\pi$ is:
+$\mathbb{E}_{\pi\sim\Pi}\left[U(S_{\pi}^{i}\cup\{i\})-U(S_{\pi}^{i})\right]=\phi_{i}(U).$
+  1. 1.
+    If $\phi_{i}(U)\geq 0$, by applying the first inequality in C.1: $P[X\geq 0]\geq\frac{\mathbb{E}[X]}{M}.$ , we get:
+    $P_{\pi\sim\Pi}\left[U(S_{\pi}^{i}\cup\{i\})\geq U(S_{\pi}^{i})\right]\geq\frac{\phi_{i}(U)}{M}.$
+  2. 2.
+    If $\phi_{\text{shap}}(i;U)\leq 0$, applying the second inequality in C.1: $P[X\geq 0]\leq 1+\frac{\mathbb{E}[X]}{m}.$ , we get:
+    $P_{\pi\sim\Pi}\left[U(S_{\pi}^{i}\cup\{i\})\geq U(S_{\pi}^{i})\right]\leq 1+\frac{\phi_{i}(U)}{m}.$